### v1

# Geometry and cohomology of compactified Deligne–Lusztig varieties

Abstract.
For connected reductive groups, we prove that the cohomology of the structure sheaf and the canonical sheaf for compactified Deligne–Lusztig varieties associated to a tuple $(w_{1},\ldots,w_{r})$ of Weyl group elements depends only on the $F$-support of $(w_{1},\ldots,w_{r})$.
Key words and phrases:
Deligne–Lusztig variety; Finite groups of Lie type; Coherent cohomology
2020 Mathematics Subject Classification:
Primary: 20G40, Secondary: 14G17
1. Introduction
1.1. Background
Let $q$ be a power of a prime $p$ and let $k$ be an algebraic closure of the finite field $\mathbb{F}_{q}$. Let $G_{0}$ be a connected reductive $\mathbb{F}_{q}$-group and $G$ its base change to $k$. Denote by $F:G\to G$ the corresponding $q$-Frobenius. The group of fixed points $G^{F}:=G(\mathbb{F}_{q})$ of $F$ is also known as a finite group of Lie type in the literature.
Let $T_{0}\subset B_{0}$ be a maximal torus and a Borel subgroup of $G_{0}$ and $T\subset B$ their base change to $k$. Denote by $W:=N(T)/T$ the corresponding Weyl group and let $S\subseteq W$ be the set of simple reflections. For any $w\in W$, the associated Deligne–Lusztig variety is denoted by $X(w)$, and there is a smooth compactification $\overline{X}(w)$ associated to each reduced expression of $w$. In [DL], Deligne and Lusztig considered the $\ell$-adic cohomology of $X(w)$ and Galois coverings of $X(w)$ to study the irreducible $G^{F}$-representations over fields of characteristic zero.
The de Rham and mod $p$ cohomology of $X(w)$ and $\overline{X}(w)$ are less understood than their $\ell$-adic cohomology. However, when $G={\rm GL}_{2}$ and $w$ is the Coxeter element, the de Rham cohomology of $\overline{X}(w)$ is not mysterious because $\overline{X}(w)$ is isomorphic to the projective line.
The simplicity in the situation of $\overline{X}(w)$ for $G={\rm GL}_{2}$ does not carry over to higher dimensions. The smooth compactification $\overline{X}(w)$ is constructed in loc. cit. as the transversal intersection between the graph of the Frobenius and the desingularisation of the Schubert variety associated to $w$ constructed by Demazure and Hansen. Unlike the desingularisation of Schubert varieties, $\overline{X}(w)$ is not naturally an iterated $\mathbb{P}^{1}$-fibration. Thus the cohomology groups of $\overline{X}(w)$ cannot be computed in the same way.
In [Wan], we described the cohomology of the structure sheaf of $\overline{X}(w)$ for any $w\in W$ in the case of $G={\rm GL}_{n}$. Our main theorem showed that the cohomology of the structure sheaf of $\overline{X}(w)$ is always isomorphic to that of $\overline{X}(c)$, where $c$ is a Coxeter element in a parabolic subgroup of $W$ with ${\text{supp}}(c)={\text{supp}}(w):=\{s\in S|s\leq w\}$. As a consequence of our main theorem, we were able to describe the compactly supported integral $p$-adic cohomology of any Deligne–Lustig variety $X(w)$ for ${\rm GL}_{n}$.
1.2. Main results
The aim of this paper is to extend the main theorem of [Wan] to all connected reductive groups $G_{0}$.
In this general setting, we encounter a special class of elements in $W$ which enjoy nice properties [GKP], [HN]. Namely, they are the minimal length elements in the $F$-conjugacy classes of $W$ (cf. Section 4.2). In any given $F$-conjugacy class $C$, we denote by $C_{\rm min}$ the subset of minimal length elements. For example, if $G$ is ${\rm GL}_{n}$ or an inner form of ${\rm GL}_{n}$, such elements are precisely the Coxeter elements in parabolic subgroups of $W$.
We show that the study of the cohomology of the structure sheaf and canonical sheaf of $\overline{X}(w)$ for any $w\in W$ can always be reduced to that of some minimal length element in an $F$-conjugacy class.
Theorem 1.1.
Let $w\in W$. There exists an $F$-conjugacy class $C$ in $W$ such that for any $x\in C_{\rm min}$, there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}(w)},\mathcal{O}_{{\overline{X}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}(x)},\mathcal{O}_{{\overline{X}(x)}}\Big{)}$
and
$H^{k}\Big{(}{\overline{X}(w)},\omega_{{\overline{X}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}(x)},\omega_{{\overline{X}(x)}}\Big{)}$
for all $k\geq 0$. Moreover, there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}(x)},\mathcal{F}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}_{L_{I}}(x),\left.\mathcal{F}\right|_{\overline{X}_{L_{I}}(x)}\Big{)},$
for all $k\geq 0$, where $I$ is the $F$-support of $x$, $P_{I}$ and $L_{I}$ are the associated standard parabolic and Levi subgroup, and $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf) of ${\overline{X}(x)}$.
One should note that the $F$-conjugacy class $C$ in the above theorem may not contain $w$. However, for any given $w$, one can determine explicitly the minimal length elements in this $F$-conjugacy class by following our recipe given in the proof of 5.1.
Smooth compactifications of $X(w)$ given by different reduced expressions of $w$ are not necessarily isomorphic. To avoid this ambiguity, we work with Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ associated to a tuple of Weyl group elements cf. 2.2. These notions also make the construction of partial compactifications of Deligne–Lusztig varieties more straightforward and uniform. One can always recover statements for classical Deligne–Lusztig varieties by setting $r=1$.
In this generality, we give a connectedness criterion for compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, generalising a result of Bonnafé and Rouquier [BR]. We also show that they have $F$-rational singularities and thus are normal, Cohen–Macaulay, and pseudo-rational. Furthermore, we provide a list of partial compactifications $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ of Deligne–Lusztig varieties which are smooth.
The structure of the paper is as follows. We fix our notations for Deligne–Lusztig varieties and compactified Deligne–Lusztig varieties and describe their geometric properties in Section 2. In Section 3, we prove that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth when each $w_{1},\ldots,w_{r}$ comes from a braid relation in the Weyl group $W$, and we establish relations between cohomology groups for the structure sheaf and canonical sheaf on the associated compactified Deligne–Lusztig varieties.
Section 4 contains the constructions of morphisms which are $\mathbb{P}^{1}$-fibrations between certain compactified Deligne–Lusztig varieties. They are associated with $F$-cyclic conjugations in the Weyl group. In 4.6, we relate the cohomology of the structure sheaf (resp. canonical sheaf) between these compactified Deligne–Lusztig varieties. Furthermore, we deduce that the cohomology of the structure sheaf (resp. canonical sheaf) for $\overline{X}(w)$ and $\overline{X}(F_{1}^{m}(w))$ are isomorphic for all $m\geq 0$, where $F_{1}$ is a partial Frobenius morphism (2.5). The key observation in this section is that the partial Frobenius morphism $F_{1}$ induces an automorphism of the Weyl group, even though it has no well-defined inverse as a morphism of schemes.
In Section 5, we begin with proving the statement in 5.1 under the assumption that $F(S)=S$. The remaining cases come from Suzuki and Ree groups, which are treated in Section 5.2. Finally, the theorem is proved in full generality in Corollary 5.3.
Acknowledgements
The author thanks Jochen Heinloth and Georg Linden for their comments on an earlier draft of this paper. Thanks to Ulrich Görtz for helpful conversations.
Parts of the work are carried out when the author was a member of the trimester program "The Arithmetic of the Langlands Program" at the Hausdorff Research Institute for Mathematics in Bonn and when the author was a member of the Research Training Group 2553 in Essen, both funded by the DFG (Deutsche Forschungsgemeinschaft).
2. Compactified Deligne–Lusztig varieties
We begin with fixing notations related to Deligne–Lusztig varieties associated to a tuple of Weyl group elements. By adapting classical proofs, we give criterions for compactified Deligne–Lusztig varieties to be connected and smooth, respectively.
2.1. Definitions and basic properties
Let $G$, $B$, $T$, $F:G\to G$, $W$, $S$ be as in Section 1.1. Denote by $\ell:W\to\mathbb{N}\cup\{0\}$ the Bruhat length and $\leq$ the Bruhat order on $W$.
Recall that by the Bruhat decomposition, the diagonal $G$-action on the product $G/B\times G/B$ yields a bijection between the Weyl group $W$ and $G\backslash\left(G/B\times G/B\right)$. The orbit corresponding to $w\in W$ is denoted by ${O(w)}$. It can be regarded as a fibre bundle over the flag variety $G/B$ with fibres isomorphic to the Schubert cell $BwB/B$.
Similarly, for a given $r$-tuple $(w_{1},\ldots,w_{r})\in W^{r}$, one can construct an iterated fibre bundle over $G/B$ by taking fibre products of the orbits $O(w_{i})$, $i=1,\ldots,r$, which we denote by
$O(w_{1},\ldots,w_{r}):=O(w_{1})\times_{G/B}\cdots\times_{G/B}O(w_{r}).$
We use such tuple of elements $w_{1},\ldots,w_{r}\in W$ for the definition of partial compactifications. They will in particular be useful for avoiding ambiguity in having different reduced expressions of a Weyl group element.
Definition 2.1.
Let $w_{1},\ldots,w_{r}\in W$. We define the compactified orbit associated to $(w_{1},\ldots,w_{r})$ as the $k$-scheme
$O(\underline{w_{1}},\ldots,\underline{w_{r}}):=\overline{O(w_{1})}\times_{G/B}\cdots\times_{G/B}\overline{O(w_{r})},$
where for each $i=1,\ldots,r$, $\overline{O(w_{i})}$ is the Zariski closure of $O(w_{i})$ in $G/B\times G/B$.
The compactified orbit $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ can be viewed as an iterated fibre bundle over $G/B$ with fibres isomorphic to Schubert varieties $\overline{Bw_{i}B}/B$. Thus $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is reduced, separated, normal, Cohen Macaulay $k$-schemes locally of finite type. The algebraic group $G$ acts diagonally on $O(w_{1},\ldots,w_{r})$ and $O(\underline{w_{1}},\ldots,\underline{w_{r}})$.
Whenever $(w_{1}^{\prime},\ldots,w_{r}^{\prime})$ is a tuple satisfying $w_{i}^{\prime}\leq w_{i}$ for all $i$, there is a $G$-equivariant embedding $O(w_{1}^{\prime},\ldots,w_{r}^{\prime})\hookrightarrow O(\underline{w_{1}},\ldots,\underline{w_{r}})$ cf.[DMR, Proposition 2.2.16]. Moreover, there is a stratification
$O(\underline{w_{1}},\ldots,\underline{w_{r}})=\bigcup_{w_{i}^{\prime}\leq w_{i},\text{ for all }i}O(w_{1}^{\prime},\ldots w_{r}^{\prime}).$
Denote by $F:G/B\to G/B$ the map induced by the Frobenius $F:G\to G$ and $\Gamma_{F}\subseteq G/B\times G/B$ its graph.
Definition 2.2.
Let $w_{1},\ldots,w_{r}\in W$. We define the Deligne–Lusztig variety associated to $(w_{1},\ldots,w_{r})$ as the fibre product
$X(w_{1},\ldots,w_{r}):=O(w_{1},\ldots,w_{r})\times_{G/B\times G/B}\Gamma_{F}$
with respect to the projection map $\text{pr}_{0,r}:O(w_{1},\ldots,w_{r})\to G/B\times G/B$ and the inclusion map $\Gamma_{F}\hookrightarrow G/B\times G/B$. Similarly, we define the compactified Deligne–Lusztig variety associated to $(w_{1},\ldots,w_{r})$ as
$X(\underline{w_{1}},\ldots,\underline{w_{r}}):=O(\underline{w_{1}},\ldots,\underline{w_{r}})\times_{G/B\times G/B}\Gamma_{F}.$
Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are separated $k$-schemes locally of finite type. The diagonal $G^{F}$-action on $O(w_{1},\ldots,w_{r})$ (resp. $O(\underline{w_{1}},\ldots,\underline{w_{r}})$) induces a diagonal $G^{F}$-action on $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$).
Whenever $(w_{1}^{\prime},\ldots,w_{r}^{\prime})$ is a tuple satisfying $w_{i}^{\prime}\leq w_{i}$ for all $i$, there is a $G^{F}$-equivariant embedding $X(w_{1}^{\prime},\ldots,w_{r}^{\prime})\hookrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Moreover, we have a stratification
$X(\underline{w_{1}},\ldots,\underline{w_{r}})=\bigcup_{w_{i}^{\prime}\leq w_{i},\text{ for all }i}X(w_{1}^{\prime},\ldots w_{r}^{\prime}).$
When $s_{i}\in S$ for all $i=1,\ldots,r$, the tuple $(s_{1},\ldots,s_{r})$ can be regarded as an element in the free monoid generated by $S$, which we denote by $S^{*}$. By Lemma 2.3 below, $X(\underline{s_{1}},\ldots,\underline{s_{r}})$ is smooth. To simplify the notation, given any $w=s_{1}\cdots s_{m}\in S^{*}$ with $s_{i}\in S,i=1,\ldots m$, we denote $X(s_{1},\ldots,s_{m})$ by $X^{S^{*}}(w)$ and $X(\underline{s_{1}},\ldots,\underline{s_{m}})$ by $\overline{X}^{S^{*}}(w)$. The notations $O^{S^{*}}(w)$ and $\overline{O}^{S^{*}}(w)$ for the orbits are defined analogously.
Furthermore, denote by $\alpha:S^{*}\to W$ be the natural projection morphism of monoids. Let $\preceq$ be the Bruhat order in $S^{*}$ and let $\ell^{S^{*}}:S^{*}\to\mathbb{N}\cup\{0\}$ be the Bruhat length.
Lemma 2.3.
Let $w_{1},\ldots,w_{r}\in W$. The Deligne–Lusztig variety $X(w_{1},\ldots,w_{r})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$. If for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$.
Proof.
In the proof of [DL, Lemma 9.11], Deligne and Lusztig showed that the graph $\Gamma_{F}$ of the Frobenius is transversal to any smooth $G$-equivariant scheme $Y\to G/B\times G/B$. Here let $Y=O(w_{1},\ldots,w_{r})$ and $Y=O(\underline{w_{1}},\ldots,\underline{w_{r}})$, which are $G$-equivariant schemes over $G/B\times G/B$. By construction, $O(w_{1},\ldots,w_{r})$ is an iterated fibration over $G/B$ with fibres being Schubert cells, which are affine spaces. The compactified orbit $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is an iterated fibration over $G/B$ with fibres being Schubert varieties $\overline{{B}w_{i}{B}}/{B}$. Hence the smoothness of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ follows. ∎
Remark 2.4.
In general, compactified Deligne–Lusztig varieties are $F$-rational and thus pseudo-rational, normal, and Cohen–Macaulay. We will give a proof for this in 2.13.
2.2. The Lang map
There is an alternative description of Deligne–Lusztig varieties via the Lang map. Using this description, we show that $G^{F}$ acts transitively on the set of connected components of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Denote by $G^{r}$ the $r$-fold product $G\times\cdots\times G$ over ${\rm Spec}\ k$. This is again a $k$-group scheme.
Definition 2.5.
Let $F_{1}:G^{r}\to G^{r}$ be the partial Frobenius morphism given by
$(g_{0},\ldots,g_{r-1})\longmapsto(g_{1},\ldots,g_{r-1},F(g_{0})).$
Let $\mathcal{L}_{1}:G^{r}\to G^{r}$ be the Lang map associated to $F_{1}$ given by $x\mapsto x^{-1}F_{1}(x)$ for $x\in G^{r}$. In other words, we have
$\mathcal{L}_{1}:(g_{0},\ldots,g_{r-1})\longmapsto(g_{0}^{-1}g_{1},\ldots,g_{r-2}^{-1}g_{r-1},g_{r-1}^{-1}F(g_{0})).$
Similarly, we define the partial Frobenius morphism on $\left(G/B\right)^{r}$. Since $B$ is the base change of $B_{0}$ to ${\rm Spec}\ k$, we know that $G/B$ is the base change of $G_{0}/B_{0}$ to ${\rm Spec}\ k$ and has a corresponding $q$-Frobenius map $F:G/B\to G/B$. Denote by $f_{2}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ the map given by
$(x_{0},\ldots,x_{r-2},x_{r-1})\longmapsto(x_{0},\ldots,x_{r-2},F(x_{r-1})),$
which is the identity on the first $(r-1)$-entries and Frobenius on the last entry. Denote by $f_{1}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ the map
$(x_{0},\ldots,x_{r-1})\longmapsto(x_{1},\ldots,x_{r-1},x_{0}),$
which shifts the coordinates by $1$. The composition $f_{1}^{r}$ is the identity.
Definition 2.6.
We define the partial Frobenius morphism $F_{1}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ to be the composition $f_{2}\circ f_{1}$.
The partial Frobenius morphism $F_{1}$ on $\left(G/B\right)^{r}$ is an étale morphism and the composition $F_{1}^{r}$ is the $q$-Frobenius morphism on $\left(G/B\right)^{r}$.
Lemma 2.7.
The Lang map $\mathcal{L}_{1}:G^{r}\to G^{r}$ is surjective and is a Galois cover with Galois group $G^{F}$.
Proof.
Note that $G$ is geometrically connected and smooth, so the product $G^{r}$ is connected and smooth. Observe that the composition of the partial Frobenius $F_{1}^{r}$ yields the Frobenius on $G^{r}$, so [CE, Theorem 7.1] applied to $G^{r}$ yields that $\mathcal{L}_{1}$ is surjective and étale. The kernel of $\mathcal{L}_{1}$ is $\Delta(G^{F})$, where $\Delta:G\to G^{r}$ is the diagonal embedding. Thus $\mathcal{L}_{1}$ is a finite morphism.
Since the diagonal (left) action of $G^{F}$ is a free action and induces automorphisms of the cover, and the number of automorphisms is the same as the degree of $\mathcal{L}_{1}$, we know that $\mathcal{L}_{1}$ is a Galois covering with Galois group $G^{F}$ [SP, 03SF]. ∎
Lemma 2.8.
For any $w_{1},\ldots,w_{r}\in W$, there are $G^{F}$-equivariant isomorphisms
$X(w_{1},\ldots,w_{r})\overset{\sim}{\longrightarrow}\mathcal{L}_{1}^{-1}\left({B}w_{1}{B}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}{B}w_{r}{B}\right)/({B})^{r}$
and
$X(\underline{w_{1}},\ldots,\underline{w_{r}})\overset{\sim}{\longrightarrow}\mathcal{L}_{1}^{-1}\left(\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\right)/({B})^{r}.$
Furthermore, $G^{F}$ acts transitively on the connected components of $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$). In particular, $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are equi-dimensional.
Proof.
We record the proof for $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ as the proof for $X(w_{1},\ldots,w_{r})$ is the same. In this proof, we denote $\mathcal{B}:=\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$. Observe that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is isomorphic to a closed subscheme of $\left(G/B\right)^{r}$ via the projection ${\rm pr}_{0,\ldots,r-1}:\left(G/B\right)^{r+1}\to\left(G/B\right)^{r}$ to the first $r$-entries. The diagonal right $({B})^{r}$-action on $G^{r}$ is free. The subscheme $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ of $G^{r}$ is stable under the right $({B})^{r}$-action, so we may take the quotient of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ with respect to the induced free right $({B})^{r}$-action. It follows that $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is a closed subscheme of $\left(G/B\right)^{r}$. The isomorphism above then follows from verifying the definitions cf. [DL]. With the diagonal $G^{F}$-action on $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$, the isomorphism is $G^{F}$-equivariant.
We show the second assertion for $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Since the quotient map $\pi:\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)\to\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equivariant under the left $G^{F}$-action and surjective, it induces a surjective morphism $G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\to G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Via Lemma 2.7 we know that
$G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\overset{\sim}{\longrightarrow}\mathcal{B}$
is irreducible, so $G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is irreducible and $G^{F}$ acts transitively on the set of connected components of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Therefore $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equidimensional. ∎
Remark 2.9.
The $G^{F}$-action on $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is not free. In $G/B$, the point ${B}$ already has nontrivial stabiliser $B^{F}=B(\mathbb{F}_{q})$. The point $(B,\ldots,B)$ in $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is also stabilised by $B^{F}=B(\mathbb{F}_{q})$.
2.3. Criterion for connectedness
There is a simple criterion for classical Deligne–Lusztig varieties to be connected, proved by Lusztig, Bonnafé–Rouquier [BR], Görtz [Gör], and many others. In this section we prove that $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$) is connected when the tuple $(w_{1},\ldots,w_{r})$ has full $F$-support by modifying the proof of [BR].
To begin, we fix the notations for the set of $F$-orbits in $S$ and the $F$-support for elements in $W$ and $S^{*}$.
Definition 2.10.
The $F$-orbit of $w\in W$ is the set ${{\rm Orb}_{F}(w)}:=\{F^{k}(w)|k\in\mathbb{Z}_{\geq 0}\}$. Under the assumption that $F(S)=S$, we define the set
$S_{F}:=\{{{\rm Orb}_{F}(s)}|s\in S\}$
of $F$-orbits in $S$. For $w\in W$, the $F$-support of $w$ is the set
${{\text{supp}}}_{F}(w):=\{\tilde{s}\in S_{F}|\text{there exists }t\in\tilde{s}\text{ such that }t\leq w\}.$
An element $w\in W$ is called of full $F$-support if ${{\text{supp}}}_{F}(w)=S_{F}$. For a tuple $(w_{1},\ldots w_{r})\in W^{r}$, we denote
${{\text{supp}}}_{F}(w_{1},\ldots,w_{r}):={{\text{supp}}}_{F}(w_{1})\cup\cdots\cup{{\text{supp}}}_{F}(w_{r}).$
Similarly, the $F$-support for any $v\in S^{*}$ with $v=s_{1}\cdots s_{m}$ is defined as
${{\text{supp}}}_{F}(v):={{\text{supp}}}_{F}(s_{1},\ldots,s_{m}).$
Remark 2.11.
We say that $w\in W$ is a Coxeter element when $w=s_{1}\cdots s_{r}$ is of full $F$-support and each $s_{i}\in S$ is from a distinct $F$-orbit.
Proposition 2.12.
[…]
Proof.
Take the product $w_{1}\cdots w_{r}$ in $S^{*}$ and denote it by $v$. This product depends on the choices of the reduced expressions of $w_{1},\ldots,w_{r}$, but for each such $v$, there is an isomorphism
$X^{S^{*}}(v)\overset{\sim}{\longrightarrow}X(w_{1},\ldots,w_{r})$
via projection. It also follows that $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are dense open subschemes of ${\overline{X}^{S^{*}}(v)}$. Thus it suffices to prove connectedness for ${\overline{X}^{S^{*}}(v)}$.
The proof showing that ${\overline{X}^{S^{*}}(v)}$ is connected is analogous to the proof of [BR] for classical Deligne–Lusztig varieties. Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=S_{F}$ and thus ${{\text{supp}}}_{F}(v)=S_{F}$. Let $Z$ be the connected component of ${\overline{X}^{S^{*}}(v)}$ which contains the point $({B},\ldots,{B})$. Denote the stabiliser of $Z$ in $G^{F}$ by ${\rm Stab}(Z)$. Fix $v^{\prime}$ to be of minimal length such that $v^{\prime}\preceq v$ and ${{\text{supp}}}_{F}(v^{\prime})=S_{F}$. In particular, $v^{\prime}$ has the same expression as $\alpha(v^{\prime})\in W$, which is a Coxeter element. This implies that ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible [Lus], [BR].
On the other hand, there is a $G^{F}$-equivariant embedding ${\overline{X}^{S^{*}}(v^{\prime})}\to{\overline{X}^{S^{*}}(v)}$. Let $Z^{\prime}$ be the connected component of ${\overline{X}^{S^{*}}(v^{\prime})}$ containing $({B},\ldots,{B})$. Then ${\rm Stab}(Z^{\prime})\subseteq{\rm Stab}(Z)$. However, ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible, so ${\rm Stab}(Z^{\prime})=G^{F}$. Thus ${\overline{X}^{S^{*}}(v)}$ is irreducible.
Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=I$ and so ${{\text{supp}}}_{F}(v)=I$. By [DMR, Proposition 2.3.8], we know that there is a $G^{F}$-equivariant isomorphism
${\overline{X}^{S^{*}}(v)}\overset{\sim}{\longrightarrow}G^{F}/U_{I}^{F}\times^{L_{I}^{F}}\overline{X}^{S^{*}}_{L_{I}}(v),$
where $U_{I}$ is the unipotent radical of $P_{I}$ and $L_{I}$ is the Levi subgroup such that $P_{I}=U_{I}\rtimes L_{I}$. Recall that $L_{I}$ is reductive, and $\overline{X}^{S^{*}}_{L_{I}}(v)$ is the compactified Deligne–Lusztig variety for $L_{I}$ corresponding to $v$. Consider the $G^{F}$-equivariant quotient morphism
$G^{F}/U_{I}^{F}\times^{L_{I}^{F}}\overline{X}^{S^{*}}_{L_{I}}(v)\longrightarrow G^{F}/P_{I}^{F},$
whose fibres are isomorphic to $\overline{X}^{S^{*}}_{L_{I}}(v)$. In this case, we know that $v$ has full $F$-support with respect to the Weyl group of $L_{I}$, so ${\overline{X}^{S^{*}}(v)}$ has $|G^{F}/P_{I}^{F}|$ connected components. ∎
2.4. Criterion for smoothness
In [Wan, Proposition A.6], we showed that some compactified Deligne–Lusztig varieties are strongly $F$-regular. This property generalises to arbitrary compactified Deligne–Lusztig varieties.
Proposition 2.13.
For $w_{1},\ldots,w_{r}\in W$, the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F$-regular. In particular, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is $F$-rational and thus normal and Cohen-Macaulay.
Proof.
The proof is similar to the proof of [Wan, Proposition A.6]. For each $i=1,\ldots,r$, we know that $\overline{{B}w_{i}{B}}$ is strongly $F$-regular [BT, Cor. 4.1, p. 958]. Thus the product
$\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$
is strongly $F$-regular by faithfully flat descent [HH, Theorem 5.5]. Recall that the Lang map $\mathcal{L}_{1}$ is étale, so its closed fibres are smooth and thus strongly $F$-regular cf. [Smi, Ch 3, Remark 1.6, Theorem 4.4(a)]. We may now apply [Abe, Theorem 3.6] and conclude that
$\mathcal{L}_{1}^{-1}\left(\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\right)$
is strongly $F$-regular. Finally, by Lemma 2.8 and the faithfully flat descent of strong $F$-regularity [HH, Theorem 5.5], $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F$-regular. Strongly $F$-regular rings are $F$-rational, so $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal and Cohen-Macaulay by [HH, Theorem 4.2, 6.27]. ∎
Remark 2.14.
For $w\in W$, the Zariski closure $\overline{X(w)}$ in $G/B$ is not always smooth, but for any $v\in S^{*}$ that gives a reduced expression of $w$, the projection map ${\overline{X}^{S^{*}}(v)}\to\overline{X(w)}$ is a proper birational morphism [DL, Lemma 9.11].
Since $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal by 2.13, the connectedness criterion for $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ in 2.12 applies readily for irreducibility.
Since smoothness can be verified étale locally and satisfies faithfully flat descent, the same proof as in 2.13 gives an alternative proof for the smoothness result in Lemma 2.3.
Corollary 2.15.
Let $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ be as in 2.13. If for all $i=1,\ldots,r$, the Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth. In particular, ${\overline{X}^{S^{*}}(w)}$ is smooth for all $w\in S^{*}$.
Example 2.16.
If for all $i=1,\ldots,r$, $w_{i}$ is the longest element in $W_{{\text{supp}}_{F}(w_{i})}$, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
3. Braid relations in the Weyl group and birational morphisms
3.1. Smoothness via braid relations
We consider some specific elements in the Weyl group for which the compactified Deligne–Lusztig variety is smooth. To begin, we recall the relations in the Weyl group of finite reductive groups cf. [GP].
All the relations in the group presentation of the Weyl group (as a finite Coxeter group $(W,S)$) can be written as
$(st)^{m_{st}}=1,\ s,t\in S,$ (3.1)
where $m_{st}\in\mathbb{Z}$ such that $m_{ss}=1$ and $m_{st}\geq 2$ if $s\neq t$. In the case of $s\neq t$, we call (3.1) a braid relation in the Weyl group $W$. The classification of finite groups of Lie type ([DM, Classification 4.3.6]) yields that for $s\neq t$, we have $m_{st}\in\{2,3,4,6\}$.
Remark 3.1.
Whenever $(st)^{m_{st}}=1,\ s,t\in S$, is a braid relation in $W$,
$(ts)^{m_{st}}=1$ (3.2)
also holds. This is a consequence of the relation $ss=1$ for all $s\in S$. We also have relations
$(F(s)F(t))^{m_{st}}=1\hskip 14.22636pt\text{and}\hskip 14.22636pt(F(t)F(s))^{m_{st}}=1$ (3.3)
corresponding to the braid relation (3.1). When there is no ambiguity, we refer to any equation among (3.1), (3.2) and (3.3) as a braid relation in $W$.
Definition 3.2.
Let $w\in S^{*}$. If $w=w_{1}w_{2}$, for some $w_{1},w_{2}\in S^{*}$, we say that $w_{i}$ divides $w$, $i=1,2$. Moreover, we call $w_{1}$ a prefix of $w$ and $w_{2}$ a suffix of $w$.
Definition 3.3.
Let $(st)^{m}=1$, with $\ s,t\in S$ and $s\neq t$, be a braid relation in $W$. We define
$P_{m}(s,t):=st\cdots\in S^{*}$ (3.4)
such that ${{\ell^{S^{*}}}(P_{m}(s,t))}=m$ and $P_{m}(s,t)$ is a prefix of $(st)^{m}$ in $S^{*}$. When there is no ambiguity, we also use $P_{m}(s,t)$ to denote its canonical projection $\alpha(P_{m}(s,t))\in W$.
Lemma 3.4.
Let $w_{1},\ldots,w_{r}\in W$. Suppose that for all $i\in\{1,\ldots,r\}$, we have $w_{i}=P_{m}(s,t)$ for some braid relation in $W$. Then the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
Proof.
By Lemma 2.3 (resp. Corollary 2.15), it suffices to show that for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth.
We show this by using the classification of finite groups of Lie type and the classification of finite Coxeter groups. Denote $I=\{s,t\}\subseteq S$, $W_{I}\subseteq W$ the corresponding subgroup generated by $I$, and $P_{I}\subseteq G$, $L_{I}\subseteq P_{I}$ the parabolic and Levi subgroups respectively. Note that $L_{I}$ is reductive and $P_{I}/{B}$ is isomorphic to $L_{I}/({B}\cap L_{I})$, which is a flag variety and thus is smooth. We proceed by showing that $S(P_{m}(s,t))=P_{I}/{B}$.
Since $W_{I}$ is generated by two elements, the longest element is determined by the braid relation. When $L_{I}$ is of type $A_{1}\times A_{1}$ (resp. ${{\ell^{S^{*}}}(P_{2}(s,t))}=2$), $st\in W_{I}$ is the longest element. When $L_{I}$ is of type $A_{2}$ or ${}^{2}A_{2}$ (resp. ${{\ell^{S^{*}}}(P_{3}(s,t))}=3$), $sts\in W_{I}$ is the longest element. When $L_{I}$ is of type $B_{2}$ or ${}^{2}B_{2}$ (resp. ${{\ell^{S^{*}}}(P_{4}(s,t))}=4$), $stst\in W_{I}$ is the longest element. When $L_{I}$ is of type $G_{2}$ or ${}^{2}G_{2}$ (resp. ${{\ell^{S^{*}}}(P_{6}(s,t))}=6$), $ststst\in W_{I}$ is the longest element. Therefore $S(P_{m}(s,t))=P_{I}/{B}$ for $m=2,3,4,6$. ∎
Definition 3.5.
Denote by $B^{+}$ the monoid generated by $S$ and all the relations of the form
$P_{m}(s,t)=P_{m}(t,s)$
with $s,t\in S$, $t\neq s$, such that $(st)^{m}=1$ is a braid relation in $W$.
More details about this monoid can be found in [GP, §4].
3.2. Higher direct images of the structure sheaf for birational morphisms
Definition 3.6.
For any $v\in S^{*}$ such that $v=P_{m}(s,t)$ for a braid relation (3.1, 3.2, 3.3) in $W$, we set $R_{st}(v):=P_{m}(t,s)$.
Proposition 3.7.
Let $w_{1},\ldots,w_{r}\in W$ such that their corresponding Schubert varieties $\overline{{B}w_{i}{B}}/{B}$ are smooth for all $i=1,\ldots,r$. Suppose for $j\in\{1,\ldots,r\}$, the reduced expression of $w_{j}$ is of the form $w_{j}=P_{m}(s,t)$ for a braid relation in $W$. Then for all $k\geq 0$, there are $G^{F}$-equivariant isomorphisms
$H^{k}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\mathcal{O}_{X_{2}}\right)$
and
$H^{k}\left(X_{1},\omega_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\omega_{X_{2}}\right),$
for $X_{i}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ and $i=1,2$.
Proof.
By assumption, we have $w_{j}=R_{st}(w_{j})$ in $W$, so
$X(\underline{w_{1}},\ldots,\underline{w_{r}})=X(\underline{w_{1}},\ldots,\underline{R_{st}(w_{j})},\ldots,\underline{w_{r}}).$
For $X_{i}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, it follows from the assumptions and Lemma 2.3 that $X_{i}$ is smooth and projective. Note that we have $I:={\text{supp}}_{F}(w_{1}\cdots w_{r})={\text{supp}}_{F}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})={\text{supp}}_{F}(w_{1},\ldots,w_{r})$, and by 2.12 $X_{i}$ has $|G^{F}/P_{I}^{F}|$ irreducible components.
There are natural projection maps ${\rm pr}_{0,\ell(w_{i})}:\overline{O}^{S^{*}}(w_{i})\to O(\underline{w_{i}})$ for all $i=1,\ldots r$ and ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(R_{st}(w_{j}))\to O(\underline{w_{j}})$. Similarly, the projection to the 00-th, $\ell(w_{1})$-th, $\ell(w_{1})+\ell(w_{2})$-th, . . ., and $\ell(w_{1})+\cdots+\ell(w_{r})$-th coordinate yields the vertical maps in the cartesian diagrams
[…]
and
[…]
The maps $f$ and $f^{\prime}$ are $G^{F}$-equivariant proper birational morphisms. It follows from the proof of Zariski’s main theorem ([Har, Corollary 11.4]) that $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$, for $\phi=f$ (resp. $\phi=f^{\prime}$), $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$ (resp. $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$) and $X_{2}=X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Moreover, applying [CR, Corollary 3.2.10] yields $R\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$. Hence the Leray spectral sequence associated to $\phi$ degenerates. Since $\phi$ is $G^{F}$-equivariant, the isomorphism $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$ is $G^{F}$-equivariant, thus functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
$H^{k}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\mathcal{O}_{X_{2}}\right),$
for all $k\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from an equivariant version of Serre duality [Nir, Corollary 2.10]. ∎
4. $\mathbb{P}^{1}$-fibrations and Cyclic shifting
We construct $\mathbb{P}^{1}$-fibrations between compactified Deligne–Lusztig varieties and relate their cohomology for the structure sheaf and canonical sheaf. Moreover, we use these $\mathbb{P}^{1}$-fibrations to indirectly show that the partial Frobenius on compactified Deligne–Lusztig varieties does not affect the cohomology for the structure sheaf and canonical sheaf.
4.1. $\mathbb{P}^{1}$-fibrations over Deligne–Lusztig varieties
Lemma 4.1.
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$. Then there is a $G^{F}$-equivariant morphism of $k$-schemes
$\pi_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
which is Zariski-locally a $\mathbb{P}^{1}$-bundle.
Proof.
For any $t\in S$, the projection map $O(\underline{t},\underline{t})\to O(\underline{t})$ given by $(B_{0},B_{1},B_{2})\mapsto(B_{0},B_{2})$ is a $\mathbb{P}^{1}$-bundle. Recall that it is a property of the Bruhat decomposition that if $(B_{0},B_{1})\in O(t)$ and $(B_{1},B_{2})\in O(t)$ then $(B_{0},B_{2})\in O(\underline{t})$, cf. [DL, §1.2]. By setting $t=F(s)$ and base change to $O(\underline{w_{1}},\ldots,\underline{w_{r}})$, we get again a $\mathbb{P}^{1}$-bundle
${\rm pr}_{F(s)}:O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})\longrightarrow O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}).$
The partial Frobenius map induces a morphism of $k$-schemes
$F_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)}).$
As $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})$ is a closed subscheme of $O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})$, we may compose ${\rm pr}$ and $F_{1}$. It follows from the construction that
${\rm pr}_{F(s)}(X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)}))\subseteq{\rm pr}_{0,r+1}^{-1}(\Gamma_{F}),$
where ${\rm pr}_{0,r+1}$ is the projection map
${\rm pr}_{0,r+1}:O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow G/B\times G/B.$
Thus the image of the composition ${\rm pr}_{F(s)}\circ F_{1}$ is contained in $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$.
We define
$\pi_{1}:={\rm pr}_{F(s)}\circ F_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
which is $G^{F}$-equivariant by construction.
For any $(B_{0},B_{1})\in O(\underline{s})$, we have $(FB_{0},FB_{1})\in O(\underline{F(s)})$. Recall that via the projection to the first $(r+2)$-entries $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ identifies with a closed subscheme of $(G/B)^{r+1}$. Under this identification, we have an embedding
$\iota:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
given by $(B_{0},\ldots,B_{r+1})\mapsto((B_{0},B_{1}),(B_{1},\ldots,B_{r+1}))$.
Since $F(s)\in S$, for any $(B_{0},B_{1})\in O(\underline{s})$, if there exists a $B^{\prime}\in G/B$ with $(B^{\prime},FB_{1})\in O(\underline{F(s)})$, then we automatically have $(B^{\prime},FB_{0})\in O(\underline{F(s)})$. Thus $\iota$ is surjective with inverse given by $((B_{0}^{\prime},B_{1}^{\prime}),(B_{1}^{\prime},\ldots,B_{r+1}^{\prime}))\mapsto(B_{0}^{\prime},\ldots,B_{r+1}^{\prime})$. Note that here we have $(B_{r+1}^{\prime},FB_{0}^{\prime})\in O(\underline{F(s)})$. This yields that $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ is isomorphic to $O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ and that $\pi_{1}$ is the composition of the isomorphism $\iota$ with the projection
$O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
which is a Zariski locally trivial $\mathbb{P}^{1}$-bundle with a trivialising section. ∎
Remark 4.2.
The reason for the fibre product of $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ with $O(\underline{s})$ yielding a Deligne–Lusztig variety is that the orbit of pairs $(B_{0},B_{1})$ determines the orbit of $(FB_{0},FB_{1})$. This allows us to verify that the tuples of the Borel subgroups live in the fibre product of the graph of Frobenius and a product of orbits. In general, taking a fibre product of a Deligne-Lusztig variety with $O(\underline{s})$ does not necessarily yield another Deligne–Lusztig variety.
Lemma 4.3.
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$. There are $G^{F}$-equivariant étale $\mathbb{P}^{1}$-bundles
$\pi_{2}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}}),$
and
$F\circ\pi_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}).$
Proof.
Define $\pi_{2}$ by
$(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\longmapsto(B_{r+1},FB_{1},\ldots,FB_{r+1}).$
This is equivariant under the diagonal $G^{F}$-action. Since $(B_{0},B_{1})\in O(\underline{s})$, we know that $(FB_{0},FB_{1})\in O(F(\underline{s}))$. By construction we have $(B_{r+1},FB_{0})\in O(F(\underline{s}))$, so $(B_{r+1},FB_{1})\in O(F(\underline{s}))$ [DL, §1.2].
Let $Z$ be given by the fibre product
[…]
where $F=F_{1}^{r+1}$ is the Frobenius map. Points in $Z$ are denoted as tuples
$((B_{0},B_{1},\ldots,B_{r+1},FB_{0}),(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})),$
where $(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$ and
$(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ such that $FB_{0}^{\prime}=B_{r+1}$, $FB_{1}^{\prime}=FB_{1},\ldots,FB_{r}^{\prime}=FB_{r},F^{2}B_{0}^{\prime}=FB_{r+1}$. Here the only condition for $B_{0}$ is that $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$.
The morphism $d_{1}$ also fits into another commutative diagram
[…]
where ${\rm pr}_{1}:O(\underline{s})\to G/B$ is given by $(B_{0},B_{1})\mapsto B_{1}$, and $r_{2}$ is given by
$((B_{0},B_{1},\ldots,B_{r+1},FB_{0}),(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime}))\longmapsto(B_{0},B_{0}^{\prime}).$
The diagram (4.2) yields an injective morphism
$\iota:Z\longrightarrow O(\underline{s})\times_{G/B}X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})}).$
On the other hand, given $(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ and arbitrary $B_{0}\in G/B$ with $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$, we have $(B_{0},B_{1}^{\prime})\in O(\underline{F^{-1}(s)})$ and $(FB_{0},FB_{0}^{\prime})\in O(\underline{s})$. This gives
$((B_{0},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime},FB_{0}),(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime}))\in Z,$
and that $\iota$ is surjective. Hence $\iota$ is an isomorphism and $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle. Let $\{U_{\alpha}\}_{\alpha}$ be a Zariksi covering of $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ over which $d_{1}$ trivialises. Since $F:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ is surjective and étale, $\{F(U_{\alpha})\}_{\alpha}$ is an étale covering of $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ over which $\pi_{2}$ trivialises as a $\mathbb{P}^{1}$-bundle.
Let $\{V_{\alpha}\}_{\alpha}$ be a Zariski covering of $X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$ over which $\pi_{1}$ trivialises. Since $F:X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\to X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ is surjective and étale, $\{F(V_{\alpha})\}_{\alpha}$ is an étale covering of $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ over which $F\circ\pi_{1}$ trivialises as a $\mathbb{P}^{1}$-bundle.
Remark 4.4.
As morphisms for $k$-schemes, the absolute Frobenius may not have a well-defined inverse, but the induced morphism $F:W\to W$ is an automorphism of the Weyl group. This is one of the key facts used in Lemma 4.3. Our constructions are analogous to the geometric constructions for classical Deligne–Lusztig varieties in the proof of [DL, Theorem 1.6].
Remark 4.5.
The maps constructed in Lemma 4.1 and Lemma 4.3 fit into a commutative diagram
[…]
Although there is a morphism $F_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})\to X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$, the morphism $F_{1}$ does not directly yield a clear relation between the cohomology of the structure sheaf for $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$. The reason is that $F_{1}$ has degree divisible by $p$, which implies that $F_{1,*}\mathcal{O}_{X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})}$ is not isomorphic to $\mathcal{O}_{X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})}$.
However, the étale $\mathbb{P}^{1}$-bundles $\pi_{1}$ and $\pi_{2}$ are projective and flat with geometrically connected and geometrically reduced fibres, so we could establish a clear relation between the cohomology of the structure sheaves of the source and the target under pushforward. This leads to the result of the following proposition that relates the cohomology of the structure sheaf (resp. canonical sheaf) between $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$.
Proposition 4.6.
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$. There are $G^{F}$-equivariant isomorphisms for all $k\geq 0$,
$H^{k}\big{(}X_{1},\mathcal{O}_{X_{1}}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}X_{2},\mathcal{O}_{X_{2}}\big{)}$
and
$H^{k}\big{(}X_{1},\omega_{X_{1}}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}X_{2},\omega_{X_{2}}\big{)},$
where for $i=1,2$, $X_{i}=X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$, $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$, $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$, and $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$.
Proof.
We continue to use the notations of Lemma 4.3. By Lemma 2.8, 2.13, and 2.12, the schemes $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ are normal, reduced, and equidimensional with $|G^{F}/P_{I}^{F}|$ irreducible components, where $I={\text{supp}}_{F}(s,w_{1},\ldots,w_{r})$.
From the proof of Lemma 4.3 we know that
$d_{1}:Z\to X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$
is a Zariski $\mathbb{P}^{1}$-bundle. It follows that all fibres of $d_{1}$ are isomorphic to $\mathbb{P}^{1}$ and that $d_{1}$ preserves connected components. Thus we may assume without loss of generality that $Z$ and $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ are irreducible. Since $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ is normal and all fibres of $d_{1}$ are geometrically reduced, a criterion for flatness [GW, Theorem 14.131] implies that $d_{1}$ is flat and thus $\pi_{2}$ is flat by base change.
Similarly, we have a Zariski $\mathbb{P}^{1}$-bundle
$\pi_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s}),$
with $X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$ normal, so $\pi_{1}$ is flat by loc. cit. Hence the composition
$F\circ\pi_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s})$
is flat.
Denote $f:=\pi_{2}$ (resp. $F\circ\pi_{1}$), $X_{2}:=X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s})$) and $X_{1}:=X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$. The rest of the proof is similar to [Wan, Proposition 4.4]. Consider the Leray spectral sequence for $\mathcal{O}_{X_{2}}$:
$E_{2}^{i,j}=H^{i}\left(X_{2},R^{j}f_{*}\mathcal{O}_{X_{1}}\right)\implies H^{i+j}\left(X_{1},\mathcal{O}_{X_{1}}\right).$
Since $f$ is proper, and all fibres of $f$ are isomorphic to $\mathbb{P}^{1}_{k}$, we have $f_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$ [SP, Lemma 0E0L]. This isomorphism is $G^{F}$-equivariant since $f$ is $G^{F}$-equivariant.
For all $x\in X_{2}$, we have $X_{1}\times_{X_{2}}x\cong\mathbb{P}^{1}_{k}$, so $H^{j}(X_{1}\times_{X_{2}}x,\mathcal{O}_{X_{1}\times_{X_{2}}x})=0$ for all $j>0$. Hence the stalk of $R^{j}f_{*}(\mathcal{O}_{X_{1}})$ at $x$ is 00 for all $x\in X_{2}$ and $j>0$ by Grauert’s theorem cf. [Vak, Theorem 25.1.6]. We may conclude that $R^{j}f_{*}\mathcal{O}_{X_{1}}=0$ for all $j>0$. Therefore the Leray spectral sequence degenerates and we have isomorphisms
$H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right),$
for all $i\geq 0$, which are $G^{F}$-equivariant by functoriality of the Leray spectral sequence.
The same argument applies for the Zariski locally trivial $\mathbb{P}^{1}$-bundle
$\pi_{1}^{\prime}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}).$
The $G^{F}$-equivariant isomorphisms for the cohomology of $\omega_{X_{1}}$ and $\omega_{X_{2}}$ follow from Serre duality for Deligne–Mumford stacks [Nir, Corollary 2.10].
4.2. $F$-conjugacy classes and $F$-cyclic shifts
Two elements $w,w^{\prime}\in W$ are $F$-conjugate if $w^{\prime}=v^{-1}wF(v)$ for some $v\in W$. The $F$-conjugacy classes in $W$ give a partition of $W$. An element $w\in W$ is of minimal length if it has the smallest Bruhat length in its $F$-conjugacy class.
Minimal length elements were studied by Geck, Kim, and Pfeiffer [GKP], as well as He and Nie [HN]. Under the assumption $F(S)=S$, these two papers showed that for any $w\in W$, there exists a minimal length element $w^{\prime}$ in the $F$-conjugacy class of $w$ such that $w\to w^{\prime}$ (cf. 4.7). In combination with 4.6, this result gives an inductive approach for studying cohomology groups of Deligne–Lusztig varieties. Let us recall the definition of $w\to w^{\prime}$.
Definition 4.7.
We write $w\to w^{\prime}$ if there exist $w_{1},\ldots,w_{r}\in W$ and $s_{1},\ldots,s_{r-1}\in S$ such that $w_{1}=w,w_{r}=w^{\prime}$, $w_{i}=s_{i-1}w_{i-1}F(s_{i-1})$, and $\ell(w_{i})\leq\ell(w_{i-1})$ for all $i=2,\ldots,r$. We say that $w$ and $w^{\prime}$ are conjugate by $F$-cyclic shifts and write $w\approx w^{\prime}$ if there exist $w_{1},\ldots,w_{r}\in W$ and $s_{1},\ldots,s_{r-1}\in S$ such that $w_{1}=w,w_{r}=w^{\prime}$, $w_{i}=s_{i-1}w_{i-1}F(s_{i-1})$, and $\ell(w_{i})=\ell(w_{i-1})$ for all $i=2,\ldots,r$.
Remark 4.8.
Let $w,w^{\prime}\in W$ such that $w\approx w^{\prime}$. If $v\in S^{*}$ gives a fixed expression for $w$, then $w\approx w^{\prime}$ guarantees that there exists $w_{1},w_{2}\in S^{*}$ such that $v=w_{1}w_{2}$ and $w_{2}F(w_{1})\in S^{*}$ gives a reduced expression for $w^{\prime}$.
The assumption $F(S)=S$ only excludes the cases when $G$ is of type ${}^{2}B_{2}$, ${}^{2}F_{4}$, or ${}^{2}G_{2}$, which will be treated separately in Section 5.2.
Proposition 4.9.
Assume $F(S)=S$. Let $w=w_{1}w_{2}$ and $w^{\prime}=w_{2}F(w_{1})$ for some $w_{1},w_{2}\in S^{*}$. For all $k\geq 0$, we have $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(w^{\prime})}}\Big{)}$
and
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w^{\prime})},\omega_{{\overline{X}^{S^{*}}(w^{\prime})}}\Big{)}.$
Proof.
This follows from 4.6. We have $w^{\prime}=F_{1}^{{{\ell^{S^{*}}}(w_{1})}}(w)$ and the result follows from composing the corresponding isomorphisms on cohomology ${{\ell^{S^{*}}}(w_{1})}$-times. ∎
Corollary 4.10.
Assume $F(S)=S$. Let $w\in S^{*}$ such that $w=w_{1}ssw_{2}$ for some $s\in S,w_{1},w_{2}\in S^{*}$. There are $G^{F}$-equivariant isomorphisms for all $k\geq 0$,
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))},\mathcal{O}_{{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))}}\Big{)},$ (4.3)
and
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))},\omega_{{\overline{X}^{S^{*}}(sw_{2}F(w_{1})F(s))}}\Big{)}.$ (4.4)
Moreover,
$H^{k}\big{(}{\overline{X}^{S^{*}}(w)},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(w_{2}F(w_{1})F(s))},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(sw_{2}F(w_{1}))},\mathcal{F}\big{)},$ (4.5)
for all $k\geq 0$ and $\mathcal{F}$ being the corresponding structure sheaf and canonical sheaf.
Proof.
As $sw_{2}F(w_{1})F(s)=F_{1}^{{{\ell^{S^{*}}}(w_{1}s)}}(w)$, the isomorphisms (4.3) and (4.4) follow from 4.9, and the isomorphisms (4.5) follow from composing (4.3) (resp. (4.4)) with the isomorphisms in 4.6. ∎
Example 4.11.
In the case when $w=ss$, we have the following isomorphisms on cohomology for for all $k\geq 0$ and for $\mathcal{F}$ being the structure sheaf (resp. canonical sheaf)
$H^{k}\big{(}{\overline{X}^{S^{*}}(ss)},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(sF(s))},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(F(s))},\mathcal{F}\big{)}\overset{\sim}{\longrightarrow}H^{k}\big{(}{\overline{X}^{S^{*}}(s)},\mathcal{F}\big{)}.$
5. Cohomology of compactified Deligne–Lustig varieties
5.1. Main theorem and applications
Theorem 5.1.
Assume $F(S)=S$. For each $w\in S^{*}$, there exists an $F$-conjugacy class $C$ in $W$ such that for any $x\in C_{\rm min}$ and any reduced expression of $x$, there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\mathcal{O}_{\overline{X}(x)}\Big{)}$
and
$H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\omega_{\overline{X}(x)}\Big{)},$
[…]
$H^{k}\Big{(}\overline{X}(x),\mathcal{F}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}_{L_{I}}(x),\left.\mathcal{F}\right|_{\overline{X}_{L_{I}}(x)}\Big{)},$
where $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf) of ${\overline{X}(x)}$.
Proof.
We give the proof for the structure sheaf. The case for the canonical sheaf follows from Serre duality for Deligne–Mumford stacks [Nir, Corollary 2.10].
By 3.7, we may assume ${\overline{X}^{S^{*}}(w)}$ has the same cohomology as an element in $B^{+}$ (3.5). In other words, let $u^{\prime}\in B^{+}$ be the image of $w$ under $S^{*}\twoheadrightarrow B^{+}$ and let $u\in S^{*}$ be a lift of $u^{\prime}$ with the same expression as $u^{\prime}$, then we have
$H^{k}\Big{(}{\overline{X}^{S^{*}}(u)},\mathcal{O}_{{\overline{X}^{S^{*}}(u)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}$
for all $k\geq 0$.
Using the description of the submonoid of reduced elements in $B^{+}$ [GP, Remark 4.1.1, Exercise 4.1], we know that the criterion for $u$ to come from a reduced expression in the Weyl group is that $u$ has no subexpression of the form $ss$ for all $s\in S$. It follows from Corollary 4.10 that there exists $v\in S^{*}$ such that $v$ has no subexpression of the form $ss$ for all $s\in S$ and for all $k\geq 0$,
$H^{k}\Big{(}{\overline{X}^{S^{*}}(v)},\mathcal{O}_{{\overline{X}^{S^{*}}(v)}}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}.$
In other words, we may assume $v\in S^{*}$ gives a reduced expression of $\alpha(v)$ in $W$. By [GKP, Theorem 2.6], there exists $v^{\prime}\in W$ with $\alpha(v)\to v^{\prime}$ and $\ell(v^{\prime})\leq\ell(\alpha(v))$. If $\alpha(v)$ is already the $F$-minimal length element in its conjugacy class, then there is nothing to prove.
In general, $\alpha(v)\to v^{\prime}$ gives $v_{1}^{\prime},\ldots,v_{m}^{\prime}\in W,s_{1},\ldots s_{m-1}\in S$ such that $v_{1}^{\prime}=\alpha(v),v_{m}^{\prime}=v^{\prime}$ and $v_{i}^{\prime}=s_{i-1}v_{i-1}^{\prime}F(s_{i-1})$ for all $i=2,\ldots,m$. Let $j\in\{1,\ldots,m\}$ be the smallest integer such that $\ell(v_{j})<\ell(v_{j-1})$. For all $i=1,\ldots j$, let $v_{i}\in S^{*}$ be the canonical lift of the corresponding reduced expression of $v_{i}^{\prime}$. It follows from 4.9 that ${\overline{X}^{S^{*}}(v_{i})}$ all have the same cohomology for the structure sheaf as ${\overline{X}^{S^{*}}(v)}$ for $i=2,\ldots,j-1$. On the other hand, we have $v_{j-1}=s_{j-1}v_{j}F(s_{j-1})$ in $S^{*}$. Via 4.6 we know that ${\overline{X}^{S^{*}}(v)}$ has the same cohomology as ${\overline{X}^{S^{*}}(s_{j-1}v_{j})}$ and ${\overline{X}^{S^{*}}(v_{j}F(s_{j-1}))}$. In this case, we reset $v$ to be $s_{j-1}v_{j}$. This process will only terminate and yields a minimal length element $x$ cf. loc. cit..
Any two minimal length elements $x$, $x^{\prime}$ in an $F$-conjugacy class $C$ are conjugate by $F$-cyclic shifts by [GKP, Theorem 2.6] and the corresponding cohomology of ${\overline{X}^{S^{*}}(x)}$ and ${\overline{X}^{S^{*}}(x^{\prime})}$ for the structure sheaf (resp. canonical sheaf) are isomorphic by 4.9.
For the last assertion, there are $G^{F}$-equivariant morphisms (cf. [DMR, Proposition 2.3.8])
${\overline{X}^{S^{*}}(x)}\overset{\sim}{\longrightarrow}G^{F}/U_{I}^{F}\times^{L_{I}^{F}}{\overline{X}^{S^{*}}_{L_{I}}(x)}\overset{\pi_{I}}{\longrightarrow}G^{F}/P_{I}^{F},$
where $P_{I}=\bigcup_{v\in W_{I}}{B}v{B}$, $U_{I}$ is the unipotent radical of $P_{I}$, and $L_{I}$ is the Levi subgroup of $G$ satisfying $P_{I}=L_{I}\ltimes U_{I}$. Since $G^{F}$ acts on the fibres of $\pi_{I}$ transitively, we have an $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}{\overline{X}^{S^{*}}(x)},\mathcal{O}_{{\overline{X}^{S^{*}}(x)}}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}^{S^{*}}_{L_{I}}(x),\mathcal{O}_{\overline{X}^{S^{*}}_{L_{I}}(x)}\Big{)}.$
for all $k\geq 0$. ∎
Remark 5.2.
With the same notations as above, it follows from the proof of the theorem that ${\rm supp}_{F}(x)={\rm supp}_{F}(w)$ for all $x\in C_{\rm min}$. Note that $x$ is of full $F$-support in $L_{I}$, so ${\overline{X}(x)}$ is irreducible by 2.12.
Corollary 5.3.
Assume $F(S)=S$. Let $v_{1},\ldots,v_{r}\in W$. There exists an $F$-conjugacy class $C$ in $W$ such that for any $x\in C_{\rm min}$ and any reduced expression of $x$, there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\mathcal{O}_{\overline{X}(x)}\Big{)}$
and
$H^{k}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\omega_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\overline{X}(x),\omega_{\overline{X}(x)}\Big{)},$
[…]
$H^{k}\Big{(}\overline{X}(x),\mathcal{F}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}(\mathbb{F}_{q})}^{G(\mathbb{F}_{q})}H^{k}\Big{(}\overline{X}_{L_{I}}(x),\left.\mathcal{F}\right|_{\overline{X}_{L_{I}}(x)}\Big{)},$
where $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf) of $\overline{X}(x)$.
Proof.
Fix reduced expressions for $v_{1},\ldots,v_{r}\in W$, and take lifts $v_{1}^{\prime},\ldots,v_{r}^{\prime}\in S^{*}$ corresponding to these reduced expressions. By 5.1, it suffices to show that the cohomology of $X(\underline{v_{1}},\ldots,\underline{v_{r}})$ for the structure sheaf is given by the cohomology of ${\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}$.
Via projection, there is a proper birational morphism $f:{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}\to X(\underline{v_{1}},\ldots,\underline{v_{r}})$ of projective and strongly $F$-regular schemes (2.13). Thus there is a $G^{F}$-equivariant isomorphism
$\phi_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\overset{\sim}{\longrightarrow}\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})},$
by the proof of Zariski’s main theorem in [Har, Corollary 11.4].
The definition of local rings being strongly $F$-regular implies that they are $F$-rational. Smith has shown that $F$-rational implies pseudo-rational [Smi, Theorem 3.1]. Moreover, recall that compactified Deligne–Lusztig varieties are normal and Cohen–Macaulay (2.13). By [Kov, Theorem 8.13], the higher direct image $R^{i}\phi_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}$ for $i>0$ vanish.
Functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
$H^{k}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\Big{)},$
for all $k\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from Serre duality for Deligne–Mumford stacks [Nir, Corollary 2.10].
The following proposition shows that taking the image of a compactified Deligne–Lusztig variety in a partial flag variety does not affect its cohomology for the structure sheaf and canonical sheaf. In particular, Corollary 5.3 applies.
Proposition 5.4.
Assume $F(S)=S$. Let $P$ be a parabolic subgroup of $G$ containing ${B}$ and denote $\pi_{P}:G/B\to G/P$ the projection map. Then there are $G^{F}$-equivariant isomorphisms
$H^{k}\Big{(}X(\underline{v}),\mathcal{F}_{X(\underline{v})}\Big{)}\overset{\sim}{\longrightarrow}H^{k}\Big{(}\pi_{P}(X(\underline{v})),\mathcal{F}_{\pi_{P}(X(\underline{v}))}\Big{)},$
for all $k\geq 0$, where $\mathcal{F}$ is the structure sheaf (resp. canonical sheaf).
Proof.
The map $\pi_{P}:G/B\to G/P$ is Zariski locally trivial with fibres isomorphic to $P/{B}\cong L/(L\cap{B})$, where $L$ is the Levi subgroup of $P$. The induced map $\left.\pi_{P}\right|_{X(\underline{v})}:X(\underline{v})\to\pi_{P}(X(\underline{v}))$ is proper and Zariski locally trivial with fibres isomorphic to $L/(L\cap{B})$. The Levi subgroup $L$ is a reductive group scheme over $k$ with finitely many connected components. Without loss of generality, we assume that $L$ is connected. Moreover, $L/(L\cap{B})$ is a flag variety over $k$, so its cohomology for the structure sheaf is given by Kempf’s vanishing theorem: $H^{j}(L/(L\cap{B}),\mathcal{O}_{L/(L\cap{B})})=0$ for $j>0$ and $H^{0}(L/(L\cap{B}),\mathcal{O}_{L/(L\cap{B})})=k$. Hence the proof of 4.6 applies. ∎
5.2. Suzuki and Ree groups
In the classification of finite reductive groups, the Frobenius endomorphism does not necessarily fix the length of elements of $S$ for groups of type ${}^{2}B_{2}$, ${}^{2}F_{4}$, and ${}^{2}G_{2}$.
Analogous to [DL, §11], we work with groups of type ${}^{2}B_{2}$, ${}^{2}F_{4}$ (resp. ${}^{2}G_{2}$) assuming $k$ is of characteristic $2$ (resp. characteristic $3$). In this situation, let $q$ be an odd power of $\sqrt{2}$ (resp. $\sqrt{3}$) and let $G$ be defined over $\mathbb{F}_{q^{2}}$. Let $F^{\prime}:G\to G$ be a Frobenius root (cf. [DM, §4.2]) such that $F^{\prime 2}$ is the the $q^{2}$-Frobenius on $G$. In order to make $F^{\prime}$ fix the length of the simple roots, fix a $F^{\prime}$-stable maximal torus $T^{*}\subseteq G$ and a Borel ${B}$ containing $T^{*}$.
A Frobenius root shares all the properties of the Frobenius morphism we used in our proofs. In particular, it is étale and its corresponding Lang map is surjective. A systematic referece for Frobenius roots can be found in loc. cit. Therefore, once we fix the datum above, 5.1 and Corollary 5.3 readily yield $G^{F^{\prime}}$-equivariant isomorphism of cohomology groups of compactified Deligne–Lustig varieties. Note that the group $G^{F^{\prime}}$ for type ${}^{2}B_{2}$ is known as the Suzuki group and $G^{F^{\prime}}$ is known as the Ree groups for type ${}^{2}F_{4}$ and ${}^{2}G_{2}$.