### v2

# Geometry and cohomology of compactified Deligne–Lusztig varieties

Abstract.
For connected reductive groups together with a Frobenius root $F$ that preserves the set of simple reflections, we show that for every tuple $(w_{1},\ldots,w_{r})$ of Weyl group elements there exists a minimal length element in an $F$-conjugacy class in $W$ such that the cohomology of the structure sheaf and the canonical sheaf for the associated compactified Deligne–Lusztig varieties are isomorphic respectively.
Key words and phrases:
Deligne–Lusztig variety; Finite groups of Lie type; Coherent cohomology
2020 Mathematics Subject Classification:
Primary: 20G40, Secondary: 14G17
1. Introduction
1.1. Background
Let $q$ be a power of a prime $p$ and let $k_{0}=\mathbb{F}_{q^{d}}$ for a positive integer $d$. Let $k$ be an algebraic closure of $\mathbb{F}_{p}$ containing $k_{0}$. Let $G_{0}$ be a connected reductive $k_{0}$-group with $q^{d}$-Frobenius $F_{0}:G_{0}\to G_{0}$. Let $G$ be the base change of $G_{0}$ to $k$ and let $F:G\to G$ be an endomorphism such that $F^{d}=F_{0}\otimes_{k_{0}}{\rm Id}$. The group of fixed points $G^{F}$ is known as a finite group of Lie type.
Let $T\subset B$ be an $F$-stable maximal torus and an $F$-stable Borel subgroup of $G$. Denote by $W:=N(T)/T$ the corresponding Weyl group and let $S\subseteq W$ be the set of simple reflections. For any $w\in W$, the associated Deligne–Lusztig variety is denoted by $X(w)$. Throughout the introduction, we fix a reduced expression of $w$ and denote by $\overline{X}(w)$ the corresponding smooth compactification of $X(w)$.
In [DL], Deligne and Lusztig considered the $\ell$-adic cohomology of $X(w)$ and certain Galois coverings of $X(w)$ to study the irreducible $G^{F}$-representations over fields of characteristic zero. The de Rham and mod $p$ cohomology of $X(w)$ and $\overline{X}(w)$ are less understood than their $\ell$-adic counterpart. However, for $G^{F}={\rm GL}_{2}(\mathbb{F}_{q})$ and $w$ the Coxeter element, the de Rham cohomology of $\overline{X}(w)$ is not mysterious because $\overline{X}(w)$ is isomorphic to the projective line.
The simplicity in the situation of $\overline{X}(w)$ for ${\rm GL}_{2}(\mathbb{F}_{q})$ does not carry over to higher dimensions. The smooth compactification $\overline{X}(w)$ is constructed in loc. cit. as the transversal intersection between the graph of $F$ and the desingularisation of the Schubert variety associated to $w$ constructed by Demazure and Hansen. Unlike these desingularisations of Schubert varieties, $\overline{X}(w)$ is not naturally an iterated $\mathbb{P}^{1}$-fibration. Thus the cohomology groups of $\overline{X}(w)$ cannot be computed in the same way.
In [Wang22], we described the cohomology of the structure sheaf of $\overline{X}(w)$ for any $w\in W$ in the case of ${\rm GL}_{n}(\mathbb{F}_{q})$. Our main theorem showed that the cohomology of the structure sheaf of $\overline{X}(w)$ is always isomorphic to that of $\overline{X}(c)$, where $c$ is a Coxeter element in a parabolic subgroup of $W$ with ${\text{supp}}(c)={\text{supp}}(w):=\{s\in S\ |\ s\leq w\}$. As a consequence of this theorem, we were able to describe the compactly supported integral $p$-adic cohomology of any Deligne–Lusztig variety $X(w)$ for ${\rm GL}_{n}(\mathbb{F}_{q})$.
1.2. Main results
The aim of this paper is to extend the main theorem of [Wang22] to all pairs $(G,F)$ for which $F(S)=S$. The assumption on $F$, for example, is satisfied when $G$ is split over $k_{0}$, or when $G^{F}$ is the group of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$.
If $G^{F}$ is a finite simple group of Lie type arising from a pair $(G,F)$ as in in Section 1.1, then the classification of such groups implies that $F(S)=S$ is satisfied. In fact, $F$ would be the Frobenius endomorphism (i.e. $d=1$) except for the cases for Suzuki and Ree groups, for which $d=2$ and additional assumptions on $q$ are required (Section 5.2).
In our setting, we encounter the minimal length elements in the $F$-conjugacy classes of $W$ (cf. Section 4.2), which is a special class of elements in $W$ with nice properties [GeckKimPfeifferMinlength, Theorem 2.6], [HeNieminlength, Theorem 3.1]. For example, if $G^{F}$ is ${\rm GL}_{n}(\mathbb{F}_{q})$ or the group of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$, then such elements are precisely the Coxeter elements in parabolic subgroups of $W$.
We show that the study of the cohomology of the structure sheaf and the canonical sheaf of $\overline{X}(w)$ for any $w\in W$ can always be reduced to that of a minimal length element in an $F$-conjugacy class in $W$.
Theorem 5.1.1.
Assume $F(S)=S$ and let $w\in W$. There exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
$H^{i}\left({\overline{X}(w)},\mathcal{F}_{{\overline{X}(w)}}\right)\overset{\sim}{\longrightarrow}H^{i}\left({\overline{X}(x_{0})},\mathcal{F}_{{\overline{X}(x_{0})}}\right),$
where $\mathcal{F}\in\{\mathcal{O},\omega\}$ is either the structure sheaf or the canonical sheaf.
Note that the $F$-conjugacy class $C$ in the theorem may not contain $w$. However, for any given $w$, one can determine $x_{0}$ and $C$ explicitly by following our recipe given in the proof of Theorem 5.1.1. We will also show that the cohomology groups in the theorem do not depend on the choice of reduced expressions of $w$ and $x_{0}$ and their corresponding smooth compactifications.
On the other hand, smooth compactifications of $X(w)$ given by different reduced expressions of $w$ are not necessarily isomorphic. To avoid this ambiguity, we work with compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ associated to a tuple of Weyl group elements, cf. Definition 2.1.1, [DMR, Definition 2.3.2]. This notion makes the construction of compactifications of Deligne–Lusztig varieties more straightforward and uniform. For example, setting $r=1$ yields the Zariski closure $X(\underline{w_{1}})=\overline{X(w_{1})}$ of $X(w_{1})$ in $G/B$. Setting $w_{1},\ldots,w_{r}\in S$ such that $w_{1}\cdots w_{r}\in W$ is a reduced expression yields the smooth compactification $X(\underline{w_{1}},\ldots,\underline{w_{r}})=\overline{X}(w_{1}\cdots w_{r})$ defined by Deligne and Lusztig.
In this generality, we give a connectedness criterion for compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, generalising a result of Bonnafé and Rouquier [BonnafeRouquierOntheirred]. We also show that they have $F_{p}$-rational singularities (Section 2.4) and thus are normal, Cohen–Macaulay, and pseudo-rational. Furthermore, we provide a list of compactifications $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ of Deligne–Lusztig varieties which are smooth.
The structure of the paper is as follows. We fix our notations for Deligne–Lusztig varieties and compactified Deligne–Lusztig varieties and describe their geometric properties in Section 2. In Section 3, we show that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth when each $w_{i},i=1,\ldots,r,$ comes from a braid relation in the Weyl group $W$, and we establish relations between cohomology groups for the structure sheaf and the canonical sheaf on the associated compactified Deligne–Lusztig varieties.
Section 4 contains the constructions of morphisms which are $\mathbb{P}^{1}$-fibrations between certain compactified Deligne–Lusztig varieties. They are associated with $F$-cyclic conjugations in the Weyl group. In Proposition 4.1.4, we relate the cohomology of the structure sheaf (resp. canonical sheaf) between these compactified Deligne–Lusztig varieties.
Finally, we prove Theorem 5.1.1 and extend the result to arbitrary compactified Deligne–Lusztig varieties of the form $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ in Section 5.
2. Compactified Deligne–Lusztig varieties
We begin with fixing notations related to Deligne–Lusztig varieties associated to a tuple of Weyl group elements. This construction has been treated in [DL, §9] and [DMR, §2]. We review here the definitions and properties that are relevant for our context. Moreover, by adapting classical proofs, we give criterions for compactified Deligne–Lusztig varieties to be connected and show that they are $F_{p}$-rational.
2.1. Definitions and basic properties
Let $G$, $F:G\to G$, $T\subset B$, $S\subset W$ be as in Section 1.1. In the literature, the endomorphism $F$ is referred to as a Frobenius root or a Steinberg endomorphism. Denote by $\ell:W\to\mathbb{N}\cup\{0\}$ the Bruhat length and $\leq$ the Bruhat order on $W$.
Recall that by the Bruhat decomposition, the diagonal $G$-action on the product $G/B\times G/B$ yields a bijection between the Weyl group $W$ and $G\backslash\left(G/B\times G/B\right)$. The orbit corresponding to $w\in W$ is denoted by ${O(w)}$. It can be regarded as a fibre bundle over the flag variety $G/B$ with fibres isomorphic to the Schubert cell $BwB/B$.
Similarly, for a given $r$-tuple $(w_{1},\ldots,w_{r})\in W^{r}$, one can construct an iterated fibre bundle over $G/B$ by taking fibre products of the orbits $O(w_{i})$, $i=1,\ldots,r$, which we denote by
$O(w_{1},\ldots,w_{r}):=O(w_{1})\times_{G/B}\cdots\times_{G/B}O(w_{r}).$
Moreover, we denote the fibre product of the Zariski closures $\overline{O(w_{i})}\subseteq G/B\times G/B$ of the orbits by
$O(\underline{w_{1}},\ldots,\underline{w_{r}}):=\overline{O(w_{1})}\times_{G/B}\cdots\times_{G/B}\overline{O(w_{r})}.$
We refer to them as the compactified orbit associated to $(w_{1},\ldots,w_{r})$ in the rest of the article. The algebraic group $G$ acts diagonally on $O(w_{1},\ldots,w_{r})$ and $O(\underline{w_{1}},\ldots,\underline{w_{r}})$.
Since $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is an iterated fibre bundle over $G/B$ with fibres isomorphic to Schubert varieties $\overline{Bw_{i}B}/B$, cf. [Jantzen, F.23], it is an integral, separated, normal, Cohen–Macaulay $k$-scheme locally of finite type. A proof of normality can be found in [DMR, Lemma 2.2.13]. The Cohen–Macaulay property behaves well under base change [stacks-project, Lemma 045T], so $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is Cohen–Macaulay follows from the classical result that Schubert varieties are Cohen–Macaulay (see the introduction of [Hashimoto]).
Denote by $F:G/B\to G/B$ the map induced by the endomorphism $F:G\to G$ and $\Gamma_{F}\subseteq G/B\times G/B$ its graph. Since $F^{d}$ is a Frobenius endomorphism, we know that $F$ is surjective, finite ([DigneMichelBook, Proposition 8.1.2]), separated ([GoerWedhAlgGeom1, Proposition 9.13]), and purely inseparable ([EGA1, (3.7.6)]). Moreover, the induced map $F:W\to W$ is a group automorphism [DigneMichelBook, Lemma 4.2.21].
Definition 2.1.1.
Let $w_{1},\ldots,w_{r}\in W$. We define the Deligne–Lusztig variety associated to $(w_{1},\ldots,w_{r})$ as the fibre product
$X(w_{1},\ldots,w_{r}):=O(w_{1},\ldots,w_{r})\times_{G/B\times G/B}\Gamma_{F}$
with respect to the projection map $\text{pr}_{0,r}:O(w_{1},\ldots,w_{r})\to G/B\times G/B$ and the inclusion map $\Gamma_{F}\hookrightarrow G/B\times G/B$. Similarly, we define the compactified Deligne–Lusztig variety associated to $(w_{1},\ldots,w_{r})$ as
$X(\underline{w_{1}},\ldots,\underline{w_{r}}):=O(\underline{w_{1}},\ldots,\underline{w_{r}})\times_{G/B\times G/B}\Gamma_{F}.$
Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are separated $k$-schemes locally of finite type. The diagonal $G^{F}$-action on $O(w_{1},\ldots,w_{r})$ (resp. $O(\underline{w_{1}},\ldots,\underline{w_{r}})$) induces a diagonal $G^{F}$-action on $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$).
For any $w_{1},\ldots,w_{r}\in W$, the $k$-scheme $X(w_{1},\ldots,w_{r})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$. If for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$, cf. [DMR, Proposition 2.3.5, 2.3.6 (i)]. In general, compactified Deligne–Lusztig varieties have $F_{p}$-rational singularities, as we will see in Proposition 2.4.1.
Whenever $(w_{1}^{\prime},\ldots,w_{r}^{\prime})$ is a tuple satisfying $w_{i}^{\prime}\leq w_{i}$ for all $i$, there is a $G^{F}$-equivariant embedding $X(w_{1}^{\prime},\ldots,w_{r}^{\prime})\hookrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Moreover, we have a stratification
$X(\underline{w_{1}},\ldots,\underline{w_{r}})=\bigcup_{w_{i}^{\prime}\leq w_{i},\text{ for all }i}X(w_{1}^{\prime},\ldots w_{r}^{\prime}).$
When $s_{i}\in S$ for all $i=1,\ldots,r$, the tuple $(s_{1},\ldots,s_{r})$ can be regarded as an element in the free monoid generated by $S$, which we denote by $S^{*}$. For any $s\in S$, the Zariski closure $\overline{BsB}=B\langle s\rangle B$ is the parabolic subgroup corresponding to the subgroup $\langle s\rangle\subset W$, cf. [DMR, Corollary 2.2.10], so the compactified Deligne–Lusztig variety $X(\underline{s_{1}},\ldots,\underline{s_{r}})$ is always smooth. To simplify the notations, given any $w=s_{1}\cdots s_{m}\in S^{*}$ with $s_{i}\in S,i=1,\ldots m$, we denote $X(s_{1},\ldots,s_{m})$ by $X^{S^{*}}(w)$ and $X(\underline{s_{1}},\ldots,\underline{s_{m}})$ by $\overline{X}^{S^{*}}(w)$. The notations $O^{S^{*}}(w)$ and $\overline{O}^{S^{*}}(w)$ for the orbits are defined analogously.
Furthermore, denote by $\alpha:S^{*}\to W$ be the natural projection morphism of monoids. Let $\preceq$ be the Bruhat order in $S^{*}$ and let $\ell^{S^{*}}:S^{*}\to\mathbb{N}\cup\{0\}$ be the Bruhat length.
2.2. The Lang map
We consider an alternative description of Deligne–Lusztig varieties via the Lang map. Using this description, we show that $G^{F}$ acts transitively on the set of connected components of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Denote by $G^{r}$ the $r$-fold product $G\times\cdots\times G$ over ${\rm Spec}\ k$. This is again a $k$-group scheme.
We begin with fixing notations. Let $F_{1}:G^{r}\to G^{r}$ be the endomorphism given by
$(g_{0},\ldots,g_{r-1})\longmapsto(g_{1},\ldots,g_{r-1},F(g_{0})).$
Note that $F_{1}$ can be viewed as a Frobenius root for the algebraic group $G^{r}$.
Let $\mathcal{L}_{1}:G^{r}\to G^{r}$ be the Lang map associated to $F_{1}$ given by $x\mapsto x^{-1}F_{1}(x)$ for $x\in G^{r}$. In other words, we have
$\mathcal{L}_{1}:(g_{0},\ldots,g_{r-1})\longmapsto(g_{0}^{-1}g_{1},\ldots,g_{r-2}^{-1}g_{r-1},g_{r-1}^{-1}F(g_{0})).$
Lemma 2.2.1.
For any $w_{1},\ldots,w_{r}\in W$, there are $G^{F}$-equivariant isomorphisms
$X(w_{1},\ldots,w_{r})\overset{\sim}{\longrightarrow}\mathcal{L}_{1}^{-1}\left({B}w_{1}{B}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}{B}w_{r}{B}\right)/({B})^{r}$
and
$X(\underline{w_{1}},\ldots,\underline{w_{r}})\overset{\sim}{\longrightarrow}\mathcal{L}_{1}^{-1}\left(\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\right)/({B})^{r}.$
Furthermore, $G^{F}$ acts transitively on the irreducible components of $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$). In particular, $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are equidimensional.
Proof.
We record the proof for $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ as the proof for $X(w_{1},\ldots,w_{r})$ is the same. In this proof, we denote $\mathcal{B}:=\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$.
By Lang’s theorem as formulated in [CabanEngue, Theorem 7.1], we deduce that $\mathcal{L}_{1}:G^{r}\to G^{r}$ is surjective and is a Galois cover with Galois group $G^{F}$. The subscheme $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ of $G^{r}$ is stable under the free $({B})^{r}$-action from the right, so we may take the quotient of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ with respect to the induced free right $({B})^{r}$-action. It follows that $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is a closed subscheme of $\left(G/B\right)^{r}$. Observe that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is also isomorphic to a closed subscheme of $\left(G/B\right)^{r}$ via projection to the first $r$-entries. The isomorphisms in this lemma then follows from verifying the definitions, cf. [DL, §1.11].
Moreover,
$G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\overset{\sim}{\longrightarrow}\mathcal{B}$
is irreducible. Since the quotient map $\pi:\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)\to\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equivariant under the left $G^{F}$-action and surjective, it induces a surjective morphism $G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\to G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Thus $G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is irreducible and $G^{F}$ acts transitively on the set of irreducible components of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$.
Remark 2.2.2.
The $G^{F}$-action on $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is not free. In $G/B$, the point ${B}$ already has nontrivial stabiliser $B^{F}=B(\mathbb{F}_{q})$. The point $(B,\ldots,B)$ in $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is also stabilised by $B^{F}=B(\mathbb{F}_{q})$.
2.3. Criterion for connectedness
There is a simple criterion for classical Deligne–Lusztig varieties to be connected, proved by Lusztig, Bonnafé–Rouquier [BonnafeRouquierOntheirred], Görtz [GoertzOntheconn], and many others. In this section we prove that $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$) is connected when the tuple $(w_{1},\ldots,w_{r})$ has full $F$-support by modifying the proof of [BonnafeRouquierOntheirred].
To begin, we fix the notations for the set of $F$-orbits in $S$ and the $F$-support for elements in $W$ and $S^{*}$.
Definition 2.3.1.
The $F$-orbit of $w\in W$ is the set ${{\rm Orb}_{F}(w)}:=\{F^{n}(w)|n\in\mathbb{Z}_{\geq 0}\}$. Under the assumption that $F(S)=S$, we define the set
$S_{F}:=\{{{\rm Orb}_{F}(s)}|s\in S\}$
of $F$-orbits in $S$. For $w\in W$, the $F$-support of $w$ is the set
${{\text{supp}}}_{F}(w):=\{\tilde{s}\in S_{F}|\text{there exists }t\in\tilde{s}\text{ such that }t\leq w\}.$
An element $w\in W$ is called of full $F$-support if ${{\text{supp}}}_{F}(w)=S_{F}$.
For a tuple $(w_{1},\ldots w_{r})\in W^{r}$, we denote
${{\text{supp}}}_{F}(w_{1},\ldots,w_{r}):={{\text{supp}}}_{F}(w_{1})\cup\cdots\cup{{\text{supp}}}_{F}(w_{r}).$
Similarly, the $F$-support for any $v\in S^{*}$ with $v=s_{1}\cdots s_{m}$ is defined as
${{\text{supp}}}_{F}(v):={{\text{supp}}}_{F}(s_{1},\ldots,s_{m}).$
Remark 2.3.2.
We say that $w\in W$ is a Coxeter element when $w=s_{1}\cdots s_{r}$ is of full $F$-support and each $s_{i}\in S$ is from a distinct $F$-orbit.
Proposition 2.3.3.
Let $w_{1},\ldots,w_{r}\in W$ and denote $I={{\text{supp}}}_{F}(w_{1},\ldots,w_{r})$. The Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are irreducible if and only if $I=S_{F}$. In general, they have $|G^{F}/P_{I}^{F}|$ connected components, where $W_{I}$ is the $F$-stable parabolic subgroup of $W$ generated by $\{s\in W|{{\rm Orb}_{F}(s)}\in I\}$ and $P_{I}={B}W_{I}{B}$.
Proof.
For certain reduced expressions of $w_{1},\ldots,w_{r}$, take the product $w_{1}\cdots w_{r}$ in $S^{*}$ and denote it by $v$. This product depends on the choices of the reduced expressions of $w_{1},\ldots,w_{r}$, but for each such $v$, there is an isomorphism
$X^{S^{*}}(v)\overset{\sim}{\longrightarrow}X(w_{1},\ldots,w_{r})$
via projection, cf. [DMR, §2.3.1]. It also follows that $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are dense open subschemes of ${\overline{X}^{S^{*}}(v)}$. Thus it suffices to prove that ${\overline{X}^{S^{*}}(v)}$ is connected if and only if $v$ has full $F$-support.
The proof showing that ${\overline{X}^{S^{*}}(v)}$ is connected is analogous to the proof of [BonnafeRouquierOntheirred] for classical Deligne–Lusztig varieties. Suppose $I=S_{F}$ and thus ${{\text{supp}}}_{F}(v)=S_{F}$. Let $Z$ be the connected component of ${\overline{X}^{S^{*}}(v)}$ which contains the point $({B},\ldots,{B})$. Denote the stabiliser of $Z$ in $G^{F}$ by ${\rm Stab}(Z)$. Fix $v^{\prime}$ to be of minimal length such that $v^{\prime}\preceq v$ and ${{\text{supp}}}_{F}(v^{\prime})=S_{F}$. In particular, $v^{\prime}$ is a reduced expression of $\alpha(v^{\prime})\in W$, which is a Coxeter element. This implies that ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible [LusztigCoxeterOrbits, Proposition 4.8].
On the other hand, there is a $G^{F}$-equivariant embedding ${\overline{X}^{S^{*}}(v^{\prime})}\to{\overline{X}^{S^{*}}(v)}$. Let $Z^{\prime}$ be the connected component of ${\overline{X}^{S^{*}}(v^{\prime})}$ containing $({B},\ldots,{B})$. Then ${\rm Stab}(Z^{\prime})\subseteq{\rm Stab}(Z)$. However, ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible, so ${\rm Stab}(Z^{\prime})=G^{F}$. Thus ${\overline{X}^{S^{*}}(v)}$ is irreducible. In this case, its dense open subschemes $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are also irreducible [GoerWedhAlgGeom1, Proposition 3.24].
Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=I$ and so ${{\text{supp}}}_{F}(v)=I$. By [DMR, Proposition 2.3.8], we know that there is a $G^{F}$-equivariant isomorphism
${\overline{X}^{S^{*}}(v)}\overset{\sim}{\longrightarrow}G^{F}/U_{I}^{F}\times^{L_{I}^{F}}\overline{X}^{S^{*}}_{L_{I}}(v),$
where $U_{I}$ is the unipotent radical of $P_{I}$ and $L_{I}$ is the Levi subgroup such that $P_{I}=U_{I}\rtimes L_{I}$. Recall that $L_{I}$ is reductive, and $\overline{X}^{S^{*}}_{L_{I}}(v)$ is the compactified Deligne–Lusztig variety for $L_{I}$ corresponding to $v$. Consider the $G^{F}$-equivariant quotient morphism
$G^{F}/U_{I}^{F}\times^{L_{I}^{F}}\overline{X}^{S^{*}}_{L_{I}}(v)\longrightarrow G^{F}/P_{I}^{F},$
whose fibres are isomorphic to $\overline{X}^{S^{*}}_{L_{I}}(v)$. In this case, we know that $v$ has full $F$-support with respect to the Weyl group of $L_{I}$, so ${\overline{X}^{S^{*}}(v)}$ has $|G^{F}/P_{I}^{F}|$ connected components. ∎
2.4. $F_{p}$-regularity
We denote by $F_{p}$ the (absolute) $p$-Frobenius to distinguish our notations from the notations coming from the theory of $F_{p}$-singularities in commutative algebra and algebraic geometry.
In [Wang22, Proposition A.6], we showed that some compactified Deligne–Lusztig varieties are strongly $F_{p}$-regular. This property generalises to arbitrary compactified Deligne–Lusztig varieties. As in loc. cit., we use a similar strategy as in the proof of [Lusztigunipotentch, Lemma 4.3], which uses faithfully flat descent of the smoothness property along the Lang map.
Proposition 2.4.1.
For $w_{1},\ldots,w_{r}\in W$, the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F_{p}$-regular. In particular, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is $F_{p}$-rational and thus normal and Cohen-Macaulay.
If $\overline{Bw_{i}B}$ is smooth for all $i$, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
Proof.
When $\overline{Bw_{i}B}$ is smooth for all $i$, this is proved by faithfully flat descent of smoothness in [Lusztigunipotentch, Lemma 4.3] and [DMR, Proposition 2.3.5], cf. [GoerWedhAlgGeom1, Proposition 14.59]. Analogously, we use faithfully flat descent of strongly $F_{p}$-regularity to show that any compactified Deligne–Lusztig variety is strongly $F_{p}$-regular.
For each $i=1,\ldots,r$, we know that $\overline{{B}w_{i}{B}}$ is strongly $F_{p}$-regular [BrTh, Cor. 4.1, p. 958] and thus Cohen-Macaulay [BrTh, Cor. 4.2, p. 959]. The product
$\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$
is strongly $F_{p}$-regular by faithfully flat descent along the projection map ${\rm pr}_{i}:\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\to\overline{{B}w_{i}{B}}$ [HoHu, Theorem 5.5.b]. Recall that the Lang map $\mathcal{L}_{1}$ is étale, so its closed fibres are smooth and thus strongly $F_{p}$-regular, cf. [KSmith, Ch 3, Remark 1.6, Theorem 4.4(a)]. We may now apply [Aber, Theorem 3.6] and conclude that
$\mathcal{L}_{1}^{-1}\left(\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\right)$
is strongly $F_{p}$-regular. Finally, by Lemma 2.2.1 and the faithfully flat descent of strong $F_{p}$-regularity [HoHu, Theorem 5.5], $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F_{p}$-regular.
Strongly $F_{p}$-regular rings are $F_{p}$-rational, so $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal and Cohen-Macaulay by [HoHu, Theorem 4.2, 6.27]. ∎
Remark 2.4.2.
For $w\in W$, the Zariski closure $\overline{X(w)}$ in $G/B$ is not always smooth, but for any $v\in S^{*}$ that gives a reduced expression of $w$, the projection map ${\overline{X}^{S^{*}}(v)}\to\overline{X(w)}$ is a proper birational morphism [DL, Lemma 9.11].
3. Braid relations in the Weyl group and birational morphisms
3.1. Smoothness via braid relations
We consider some specific elements in the Weyl group for which the compactified Deligne–Lusztig variety is smooth. To begin, we recall the relations in the Weyl group of finite reductive groups, cf. [GeckPfeiffer].
All the relations in the group presentation of the Weyl group as a finite Coxeter group $(W,S)$ can be written as
$(st)^{m_{st}}=1,\ s,t\in S,$ (1)
where $m_{st}\in\mathbb{Z}$ such that $m_{ss}=1$ and $m_{st}\geq 2$ if $s\neq t$. In the case of $s\neq t$, we call (1) a braid relation in the Weyl group $W$. We refer to such an equation (1) as a braid relation in $W$.
Remark 3.1.1.
If $F(S)=S$, then whenever $(st)^{m_{st}}=1,\ s,t\in S$ is a braid relation in $W$, the equations
$(ts)^{m_{st}}=1,\hskip 14.22636pt\ (F(s)F(t))^{m_{st}}=1,\hskip 14.22636pt\text{and}\hskip 14.22636pt(F(t)F(s))^{m_{st}}=1$
also hold.
For $w=w_{1}w_{2}\in S^{*}$, we call $w_{1}$ a prefix of $w$ and $w_{2}$ a suffix of $w$.
Definition 3.1.2.
Let $(st)^{m}=1$, with $\ s,t\in S$ and $s\neq t$, be a braid relation in $W$. We define
$P(s,t):=st\cdots\in S^{*},$ (2)
such that ${{\ell^{S^{*}}}(P(s,t))}=m$ and $P(s,t)$ is a prefix of $(st)^{m}$ in $S^{*}$. When there is no ambiguity, we also use $P(s,t)$ to denote its canonical projection $\alpha(P(s,t))\in W$.
Lemma 3.1.3.
Let $w_{1},\ldots,w_{r}\in W$. Suppose that for all $j\in\{1,\ldots,r\}$, we have $w_{j}=P(s,t)$ for some braid relation in $W$. Then the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
Proof.
By Proposition 2.4.1, it suffices to show that for all $j=1,\ldots,r$, the associated shceme $\overline{{B}w_{j}{B}}$ is smooth.
Consider $w_{j}=P(s,t)$ with $(st)^{m}=1$. Denote $J=\{s,t\}\subseteq S$. The the corresponding subgroup $W_{J}\subseteq W$ generated by $J$ has a group presentation
$W_{J}=\langle s,t|(st)^{m}=1,s^{2}=1,t^{2}=1\rangle.$
As a consequence, $P(s,t)$ is a reduced expression of the longest element of $W_{J}$.
Let $P_{J}=BW_{J}B\subseteq G$ be the parabolic subgroup corresponding to $W_{J}$. Since $w_{j}$ is the longest element in $W_{J}$, we know that $\overline{{B}w_{j}{B}}=P_{J}$ and is smooth cf. [DigneMichelBook, Proposition 3.2.8]. ∎
3.2. Higher direct images of the structure sheaf for birational morphisms
Definition 3.2.1.
For any $v\in S^{*}$ such that $v=P(s,t)$ for a braid relation (1) in $W$, we set $R(s,t):=P(t,s)$.
Proposition 3.2.2.
Let $w_{1},\ldots,w_{r}\in W$ such that $\overline{{B}w_{j}{B}}$ is smooth for all $j=1,\ldots,r$. Suppose for $j\in\{1,\ldots,r\}$, the reduced expression of $w_{j}$ is of the form $w_{j}=P(s,t)$ for a braid relation in $W$. Then for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
$H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right)$
and
$H^{i}\left(X_{1},\omega_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\omega_{X_{2}}\right),$
for $X_{n}$ $(n\in\{1,2\})$ being one of the $k$-schemes ${\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$, and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, where $R(s,t)$ replaces $w_{j}$.
Proof.
By assumption, we have $w_{j}=P(s,t)=R(s,t)$ in $W$, so replacing $w_{j}$ by $R(s,t)$ does not affect the following compactified Deligne–Lusztig variety
$X(\underline{w_{1}},\ldots,\underline{w_{r}})=X(\underline{w_{1}},\ldots,\underline{R(s,t)},\ldots,\underline{w_{r}}).$
Replace $w_{j}$ by $R(s,t)$ in the expression $w_{1}\cdots w_{r}\in S^{*}$ and denote the resulting element by $w_{1}\cdots R(s,t)\cdots w_{r}$. For $X_{n}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, it follows from the assumptions and Lemma 3.1.3 that $X_{n}$ is smooth and projective. Note that we have $I:={\text{supp}}_{F}(w_{1}\cdots w_{r})={\text{supp}}_{F}(w_{1}\cdots R(s,t)\cdots w_{r})={\text{supp}}_{F}(w_{1},\ldots,w_{r})$, and by Proposition 2.3.3 $X_{n}$ has $|G^{F}/P_{I}^{F}|$ irreducible components.
There are natural projection maps ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(w_{j})\to O(\underline{w_{j}})$ and ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(R(s,t))\to O(\underline{w_{j}})$ for all $j=1,\ldots r$. Similarly, the projection to the 00-th, $\ell(w_{1})$-th, $\ell(w_{1})+\ell(w_{2})$-th, . . ., and $\ell(w_{1})+\cdots+\ell(w_{r})$-th coordinate yields the vertical maps in the cartesian diagrams
${{\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}}$${\overline{O}^{S^{*}}(w_{1}\cdots w_{r})}$${X(\underline{w_{1}},\ldots,\underline{w_{r}})}$${O(\underline{w_{1}},\ldots,\underline{w_{r}}),}$$\scriptstyle{f}$
and
${{\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}}$${\overline{O}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$${X(\underline{w_{1}},\ldots,\underline{w_{r}})}$${O(\underline{w_{1}},\ldots,\underline{w_{r}}).}$$\scriptstyle{f^{\prime}}$
The maps $f$ and $f^{\prime}$ are $G^{F}$-equivariant proper birational morphisms.
It follows from the proof of Zariski’s main theorem ([Hartshorne, Corollary 11.4]) that $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$, for $\phi=f$ (resp. $\phi=f^{\prime}$), $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$ (resp. $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$) and $X_{2}=X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Moreover, applying [CR11, Corollary 3.2.10] yields $R\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$. Hence the Leray spectral sequence associated to $\phi$ degenerates.
Since $\phi$ is $G^{F}$-equivariant, the isomorphism $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$ is $G^{F}$-equivariant, thus functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
$H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right),$
for all $i\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from an equivariant version of Serre duality [Nironi, Corollary 2.10]. ∎
4. $\mathbb{P}^{1}$-fibrations and $F$-conjugacy classes
We construct $\mathbb{P}^{1}$-fibrations between compactified Deligne–Lusztig varieties and relate their cohomology for the structure sheaf and the canonical sheaf. Moreover, we use these $\mathbb{P}^{1}$-fibrations to show that pairs of compactified Deligne–Lusztig varieties of the form ${\overline{X}^{S^{*}}(sv)}$ and ${\overline{X}^{S^{*}}(vF(s))}$ have isomorphic cohomology groups for the the structure sheaf and the canonical sheaf.
4.1. $\mathbb{P}^{1}$-fibrations over Deligne–Lusztig varieties
Lemma 4.1.1.
Let $w_{1},\ldots,w_{r},w_{1}^{\prime},\ldots,w_{m}^{\prime}\in W$ and $s\in S$. Then the projection map
$\phi:X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s},\underline{s},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})$
given by
$(B_{0},\ldots,B_{r},B_{r+1},B_{r+2},\ldots,B_{r+m+1},FB_{0})\mapsto(B_{0},\ldots,B_{r},B_{r+2},\ldots,B_{r+m+1},FB_{0})$ is a $G^{F}$-equivariant Zariski $\mathbb{P}^{1}$-bundle.
If $F(s)\in S$, then the $G^{F}$-equivariant morphism of $k$-schemes
$\pi:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
given by $(B_{0},\ldots,B_{r+1},FB_{0})\mapsto(B_{1},\ldots,B_{r+1},FB_{1})$, is a Zariski $\mathbb{P}^{1}$-bundle.
Proof.
For any $t\in S$, the projection map $O(\underline{t},\underline{t})\to O(\underline{t})$ given by $(B_{0},B_{1},B_{2})\mapsto(B_{0},B_{2})$ is a $\mathbb{P}^{1}$-bundle. Recall that it is a property of the Bruhat decomposition that if $(B_{0},B_{1})\in O(t)$ and $(B_{1},B_{2})\in O(t)$ then $(B_{0},B_{2})\in O(\underline{t})$, cf. [DL, §1.2].
For $t=s$, we see that $\phi$ is the base change of $O(\underline{t},\underline{t})\to O(\underline{t})$ along the projection map $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{t},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})\to O(\underline{t})$ given by $(B_{0}^{\prime},\ldots,FB_{0}^{\prime})\mapsto(B_{r}^{\prime},B_{r+1}^{\prime})$. The first claim follows.
Assume $F(s)\in S$ and consider $t=F(s)$. One can verify that $\pi$ is $G^{F}$-equivariant by its definition. Let $(B_{0},B_{1})\in O(\underline{s})$, so $(FB_{0},FB_{1})\in O(\underline{F(s)})$. Recall that via the projection to the first $(r+2)$-entries, $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ can be identified with a closed subscheme of $(G/B)^{r+2}$. Under this identification, we have an embedding
$\iota:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
given by $(B_{0},\ldots,B_{r+1})\mapsto((B_{0},B_{1}),(B_{1},\ldots,B_{r+1}))$.
Since $F(s)\in S$, for any $(B_{0},B_{1})\in O(\underline{s})$, if there exists a $B^{\prime}\in G/B$ with $(B^{\prime},FB_{1})\in O(\underline{F(s)})$, then we automatically have $(B^{\prime},FB_{0})\in O(\underline{F(s)})$. Thus $\iota$ is surjective with inverse given by $((B_{0}^{\prime},B_{1}^{\prime}),(B_{1}^{\prime},\ldots,B_{r+1}^{\prime}))\mapsto(B_{0}^{\prime},\ldots,B_{r+1}^{\prime})$. Therefore $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ is isomorphic to $O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$. It follows that $\pi$ is the composition of the isomorphism $\iota$ with the projection
$O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
which is a Zariski $\mathbb{P}^{1}$-bundle. ∎
Remark 4.1.2.
The reason for the fibre product of $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ with $O(\underline{s})$ to yield a Deligne–Lusztig variety is that the orbit of pairs $(B_{0},B_{1})$ determines the orbit of $(FB_{0},FB_{1})$. This allows us to verify that the tuples $(B_{0},\ldots,B_{r+1},FB_{0})$ live in the fibre product of the graph of $F$ and a product of orbits. In general, taking a fibre product of a Deligne-Lusztig variety with $O(\underline{s})$ does not necessarily yield another Deligne–Lusztig variety.
Lemma 4.1.3.
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$ and $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. The morphisms
$\tau_{2}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}}),$
given by $(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\longmapsto(B_{r+1},FB_{1},\ldots,FB_{r+1})$, and
$\tau_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
given by $(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\longmapsto(FB_{1},FB_{2}\ldots,F^{2}B_{1})$, are $G^{F}$-equivariant fppf $\mathbb{P}^{1}$-bundles.
Proof.
Given $(B_{0},B_{1})\in O(\underline{s})$, we know that $(FB_{0},FB_{1})\in O(F(\underline{s}))$. By construction we have $(B_{r+1},FB_{0})\in O(F(\underline{s}))$, so $(B_{r+1},FB_{1})\in O(F(\underline{s}))$ and $(FB_{r+1},F^{2}B_{1})\in O(F(\underline{s}))$ [DL, §1.2]. Thus $\tau_{1}$ and $\tau_{2}$ are well-defined. They are $G^{F}$-equivariant by construction.
Let $Z$ be the fibre product
[…]
Geometric points in $Z$ are denoted as tuples
$((B_{0},B_{1},\ldots,B_{r+1},FB_{0}),(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})),$
where $(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$ and
$(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ such that $FB_{0}^{\prime}=B_{r+1}$, $FB_{1}^{\prime}=FB_{1},\ldots,FB_{r}^{\prime}=FB_{r},F^{2}B_{0}^{\prime}=FB_{r+1}$. Here the only condition on $B_{0}$ is that $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$.
The morphism $d_{1}$ also fits into another commutative diagram
${Z}$${O(\underline{F^{-1}(s)})}$${X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})}$${G/B,}$$\scriptstyle{r_{2}}$$\scriptstyle{d_{1}}$$\scriptstyle{{\rm pr}_{1}}$$\scriptstyle{{\rm pr}_{0}}$ (4)
where ${\rm pr}_{1}:O(\underline{s})\to G/B$ is given by $(B_{0},B_{1})\mapsto B_{1}$, and $r_{2}$ is given by
$((B_{0},B_{1},\ldots,B_{r+1},FB_{0}),(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime}))\longmapsto(B_{0},B_{0}^{\prime}).$
The diagram (4) yields a morphism
$\iota:Z\longrightarrow O(\underline{F^{-1}(s)})\times_{G/B}X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})}).$
Given $(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ and arbitrary $B_{0}\in G/B$ with $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$, we have $(B_{0},B_{1}^{\prime})\in O(\underline{F^{-1}(s)})$ and $(FB_{0},FB_{0}^{\prime})\in O(\underline{s})$. This gives
$((B_{0},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime},FB_{0}),(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime}))\in Z,$
and that $\iota$ is surjective. By the Bruhat decomposition, for every $i=1,\ldots,r$, there exists $v_{i}\in W$ such that $(B_{i},B_{i}^{\prime})\in O(v_{i})$. This implies $(FB_{i},FB_{i}^{\prime})\in O(F(v_{i}))$. We know $FB_{i}=FB_{i}^{\prime}$ by assumption, so $F(v_{i})=e$ is the neutral element in $W$ and $v_{i}=e$ because $F:W\to W$ is an automorphism, cf. [DL, §1.2]. It follows that $\iota$ is also injective and an isomorphism.
Hence $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle. Since $F:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ is faithfully flat by our assumption on smoothness [GoerWedhAlgGeom1, Corollary 14.129], we deduce that $\tau_{2}$ is faithfully flat by faithfully flat descent [GoerWedhAlgGeom1, Corollary 14.12, Proposition 14.3(2)].
For each affine open (Zariski) covering $\{U_{\alpha}\}_{\alpha}$ of $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ that trivialises $d_{1}$, we can construct an fppf covering $\{\left.F\right|_{U_{\alpha}}:U_{\alpha}\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})\}_{\alpha}$ of $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ over which $\tau_{2}$ trivialises as a $\mathbb{P}^{1}$-bundle.
After taking the fibre product of $\tau_{1}$ with $F:X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\to X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$, the claim for $\tau_{1}$ can be verified by the analogous argument we used for $\tau_{2}$ above. ∎
Proposition 4.1.4.
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$ and $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
$H^{i}\big{(}X_{1},\mathcal{O}_{X_{1}}\big{)}\overset{\sim}{\longrightarrow}H^{i}\big{(}X_{2},\mathcal{O}_{X_{2}}\big{)}$
and
$H^{i}\big{(}X_{1},\omega_{X_{1}}\big{)}\overset{\sim}{\longrightarrow}H^{i}\big{(}X_{2},\omega_{X_{2}}\big{)}$
induced by $f:X_{1}\to X_{2}$, where $f$ is one of the $\mathbb{P}^{1}$-bundles $\phi,\pi,\tau_{1}$ or $\tau_{2}$ from Lemma 4.1.1 and 4.1.3.
Proof.
The proof is simpler for the Zariski $\mathbb{P}^{1}$-bundles $\phi$ and $\pi$, so we only record the proof for the fppf $\mathbb{P}^{1}$-bundles $\tau_{1}$ and $\tau_{2}$. We use the notations of Lemma 4.1.3. Since $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle, it is proper and faithfully flat. Hence $\tau_{1}$ and $\tau_{2}$ are proper and faithfully flat by faithfully flat descent [GoerWedhAlgGeom1, Proposition 14.53].
Set $f:=\tau_{2}$ (resp. $\tau_{1}$), $X_{1}:=X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$, $X_{2}:=X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$), and $X_{2}^{\prime}:=X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ (resp. $X(\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})$). Consider the Leray spectral sequence for $\mathcal{O}_{X_{1}}$:
$E_{2}^{i,j}=H^{i}\left(X_{2},R^{j}f_{*}\mathcal{O}_{X_{1}}\right)\implies H^{i+j}\left(X_{1},\mathcal{O}_{X_{1}}\right).$
Let $x\in X_{2}$ and $y\in X_{2}^{\prime}$ such that $F(y)=x$ as in diagram (3). By Lemma 4.1.3, the base change of the fibre $x\times_{X_{2}}X_{1}$ to ${\rm Spec}(\kappa(y))$ is isomorphic to $\mathbb{P}^{1}_{\kappa(y)}$. It follows from flat base change that $H^{j}(x\times_{X_{2}}X_{1},\mathcal{O}_{x\times_{X_{2}}X_{1}})=0$ for all $j\geq 0$ [GoerWedhAlgGeom2, Corollary 22.91]. We conclude by Grauert’s theorem [GoerWedhAlgGeom2, Proposition 23.142] (cf. [GoerWedhAlgGeom1, Proposition 3.35]) that $R^{j}f_{*}\mathcal{O}_{X_{1}}=0$ for all $j>0$. Hence the Leray spectral sequence degenerates.
We have an fppf covering $\mathfrak{U}=\{\gamma_{\alpha}:U_{\alpha}\to X_{2}\}_{\alpha}$ of $X_{2}$ in Lemma 4.1.3, such that the base change $f_{\alpha}:X_{1}\times_{X_{2}}U_{\alpha}\to U_{\alpha}$ of $f$ along $\gamma_{\alpha}$ is a trivial $\mathbb{P}^{1}$-bundle for all $\alpha$. Since $U_{\alpha}$ is affine and $\gamma_{\alpha}$ is an affine morphism for all $\alpha$ [GoerWedhAlgGeom1, Proposition 12.3(3)], any simplex $V$ in the Čech nerve of $\mathfrak{U}$ is affine and for any coherent sheaf $\mathcal{F}$ on $X_{2}$, we have $H^{i}(X_{2},\mathcal{F})\overset{\sim}{\longrightarrow}\check{H}^{i}(\mathfrak{U},\mathcal{F})$ for all $i\geq 0$ [Godement, Theorem II.5.4.1, Corollary]. Note that $f_{*}\mathcal{O}_{X_{1}}$ is a coherent $\mathcal{O}_{X_{2}}$-module because $f$ is proper and cohomology of coherent sheaves is independent of topology [stacks-project, 03DW].
For each simplex $V$ in the Čech nerve of $\mathfrak{U}$, we have $V\times_{X_{2}}X_{1}\overset{\sim}{\longrightarrow}V\times_{{\rm Spec}\ k}\mathbb{P}^{1}$, so $f_{*}\mathcal{O}_{X_{1}}(V)=\mathcal{O}_{V}(V)$ by flat base change [GoerWedhAlgGeom2, Proposition 22.90]. This identification is compatible with restriction of sections, so the Čech complexes on $\mathfrak{U}$ for $f_{*}\mathcal{O}_{X_{1}}$ and $\mathcal{O}_{X_{2}}$ are identical and $H^{i}(X_{2},\mathcal{O}_{X_{2}})\overset{\sim}{\longrightarrow}H^{i}(X_{2},f_{*}\mathcal{O}_{X_{1}})$ for all $i\geq 0$.
Finally, degeneration of the Leray spectral sequence yields $G^{F}$-equivariant isomorphisms
$H^{i}\left(X_{2},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right),$
for all $i\geq 0$.
The $G^{F}$-equivariant isomorphisms for the cohomology of $\omega_{X_{1}}$ and $\omega_{X_{2}}$ follow from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
Corollary 4.1.5.
Let $s\in S$ and assume $F(s)\in S$. Let $w=sv$ and $w^{\prime}=vF(s)$ be elements of $S^{*}$. There are $G^{F}$-equivariant isomorphisms for all $i\geq 0$,
$H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(w^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(w^{\prime})}}\Big{)}$
and
$H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(w^{\prime})},\omega_{{\overline{X}^{S^{*}}(w^{\prime})}}\Big{)}.$
Proof.
This follows from Proposition 4.1.4 by composing the corresponding isomorphisms of cohomology groups for $\tau_{1}$ and $\tau_{2}$. ∎
4.2. $F$-conjugacy classes in $W$
Two elements $w,w^{\prime}\in W$ are $F$-conjugate if $w^{\prime}=v^{-1}wF(v)$ for some $v\in W$. The $F$-conjugacy classes in $W$ give a partition of $W$. An element $w\in W$ is of minimal length if it has the smallest Bruhat length in its $F$-conjugacy class.
Minimal length elements were studied by Geck, Kim, and Pfeiffer [GeckKimPfeifferMinlength], as well as He and Nie [HeNieminlength]. Under the assumption $F(S)=S$, these two papers showed that for any $w\in W$, there exists a minimal length element $w^{\prime}$ in the $F$-conjugacy class of $w$ such that $w\to_{F}w^{\prime}$ (cf. Definition 4.2.1). In combination with Proposition 4.1.4, this result gives an inductive approach for studying cohomology groups of Deligne–Lusztig varieties.
Definition 4.2.1.
We write $w\to_{F}w^{\prime}$ if there exist $w_{1},\ldots,w_{r}\in W$ and $s_{1},\ldots,s_{r-1}\in S$ such that $w_{1}=w,w_{r}=w^{\prime}$, $w_{i}=s_{i-1}w_{i-1}F(s_{i-1})$, and $\ell(w_{i})\leq\ell(w_{i-1})$ for all $i=2,\ldots,r$.
Furthermore, for any pairs of minimal length elements $w$ and $w^{\prime}$ in an elliptic $F$-conjugacy class $C$ in $W$, we have $w\to_{F}w^{\prime}$ and $w^{\prime}\to_{F}w$, cf. [HeNieminlength, Corollary 4.4]. The definition of an elliptic $F$-conjugacy class, which we do not recall here, can be found in §4.1 of loc. cit.
5. Cohomology of compactified Deligne–Lusztig varieties
5.1. Main theorem and applications
Recall that the braid monoid associated to $W$ is a monoid generated by $S$ and all the relations of the form $P(s,t)=P(t,s)$ with $s,t\in S$, $t\neq s$, such that $(st)^{m}=1$ is a braid relation in $W$, cf. [GeckPfeiffer, §4]. We denote the braid monoid by $B^{+}$. There is a natural surjective morphism of monoids $\beta:S^{*}\twoheadrightarrow B^{+}$.
Theorem 5.1.1.
Assume $F(S)=S$. For each $w\in S^{*}$, there exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for any reduced expression $x\in S^{*}$ of $x_{0}$, there are $G^{F}$-equivariant isomorphisms
$H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\mathcal{O}_{{\overline{X}^{S^{*}}(x)}}\Big{)}$
and
$H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\omega_{{\overline{X}^{S^{*}}(w)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\omega_{{\overline{X}^{S^{*}}(x)}}\Big{)},$
for all $i\geq 0$.
If $C$ is moreover an elliptic conjugacy class in $W$, then the isomorphisms above hold if we replace $x_{0}$ by any other minimal length element in $C$.
Proof.
We prove the theorem for the structure sheaf. The results for the canonical sheaf follows from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
Let $w\in S^{*}$. Consider the image $\beta(w)\in B^{+}$ of $w$ under the projection $\beta:S^{*}\twoheadrightarrow B^{+}$ and let $u\in S^{*}$ be another lift of $\beta(w)$, then we have
$H^{i}\Big{(}{\overline{X}^{S^{*}}(u)},\mathcal{O}_{{\overline{X}^{S^{*}}(u)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}$ (5)
for all $i\geq 0$ by Proposition 3.2.2. Note that this proposition can be applied to any $t_{1}\ldots t_{r}\in S^{*}$ by setting $w_{j}=t_{j}$ for $j=1,\ldots r$.
Using the description of the submonoid of reduced elements in $B^{+}$ given by Matsumoto’s theorem [GeckPfeiffer, Theorem 1.2.2, Exercise 4.1], we know that the criterion for $u$ to come from a reduced expression in the Weyl group is that $u$ has no subexpression of the form $ss$ for any $s\in S$. If $u=u_{1}ssu_{2}$, we apply Lemma 4.1.1 and Proposition 4.1.4 to obtain $u_{1}su_{2}$ such that ${\overline{X}^{S^{*}}(u)}$ and ${\overline{X}^{S^{*}}(u_{1}su_{2})}$ have the same cohomology for the structure sheaf. Inductively, this yields an element $v\in S^{*}$ with ${{\ell^{S^{*}}}(v)}\leq{{\ell^{S^{*}}}(u)}$ such that $v$ has no subexpression of the form $ss$ for any $s\in S$ and for all $i\geq 0$,
$H^{i}\Big{(}{\overline{X}^{S^{*}}(v)},\mathcal{O}_{{\overline{X}^{S^{*}}(v)}}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(w)},\mathcal{O}_{{\overline{X}^{S^{*}}(w)}}\Big{)}.$
In other words, we may assume $v\in S^{*}$ gives a reduced expression of $\alpha(v)$ in $W$.
By [GeckKimPfeifferMinlength, Theorem 2.6], there exists $v^{\prime}\in W$ which is of minimal length in its $F$-conjugacy class such that $\alpha(v)\to_{F}v^{\prime}$ and $\ell(v^{\prime})\leq\ell(\alpha(v))$. Suppose $\alpha(v)$ is a minimal length element in its $F$-conjugacy class. By Matsumoto’s theorem [GeckPfeiffer, Theorem 1.2.2], there exists a lift $y\in B^{+}$ of $\alpha(v)$ such that any reduced expression $x\in S^{*}$ of $\alpha(v)$ satisfies $\beta(x)=y$. The statement of the theorem thus follows from Proposition 3.2.2.
In general, $\alpha(v)\to_{F}v^{\prime}$ gives $v_{1}^{\prime},\ldots,v_{m}^{\prime}\in W,s_{1},\ldots s_{m-1}\in S$ such that $v_{1}^{\prime}=\alpha(v),v_{m}^{\prime}=v^{\prime}$ and $v_{j}^{\prime}=s_{j-1}v_{j-1}^{\prime}F(s_{j-1})$ for all $j=2,\ldots,m$. Let $a\in\{2,\ldots,m\}$ be the smallest integer such that $\ell(v_{a}^{\prime})<\ell(v_{a-1}^{\prime})$. For all $j=1,\ldots,a$, let $v_{j}\in S^{*}$ be the canonical lift of the corresponding reduced expression of $v_{j}^{\prime}$. Up to taking different lifts of braid relations, for each $j=2,\ldots,a-1$, we have $v_{j}=s_{j-1}z_{j-1}$ (resp. $v_{j}=z_{j-1}F(s_{j-1})$) and $v_{j-1}=z_{j-1}F(s_{j-1})$ (resp. $v_{j-1}=s_{j-1}z_{j-1}$) for some $z_{j-1}\in S^{*}$ (cf. [GeckKimPfeifferMinlength, Remark 2.3]). Thus it follows from Corollary 4.1.5 that ${\overline{X}^{S^{*}}(v_{j})}$ all have the same cohomology for the structure sheaf as ${\overline{X}^{S^{*}}(v)}$ for $j=2,\ldots,a-1$.
On the other hand, we have $v_{a-1}=s_{a-1}v_{a}F(s_{a-1})$ in $S^{*}$. We know that ${\overline{X}^{S^{*}}(v)}$ has the same cohomology as ${\overline{X}^{S^{*}}(s_{a-1}v_{a})}$ and ${\overline{X}^{S^{*}}(v_{a}F(s_{a-1}))}$ by Proposition 4.1.4. In this case, we reset $v$ to be $s_{a-1}v_{a}$. This process will terminate and yields an element $x_{0}\in W$ which is of minimal length in its $F$-conjugacy class, cf. [GeckKimPfeifferMinlength, Theorem 2.6].
It remains to consider the case when $C$ is an elliptic conjugacy class in $W$. By [HeNieminlength, Corollary 4.4], we have $x\to_{F}y$ for any two minimal length elements $x,y\in C$. The corresponding cohomology of ${\overline{X}^{S^{*}}(x)}$ and ${\overline{X}^{S^{*}}(y)}$ for the structure sheaf are isomorphic by Corollary 4.1.5. ∎
Corollary 5.1.2.
Assume $F(S)=S$. Let $v_{1},\ldots,v_{r}\in W$. There exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for any reduced expression $x\in S^{*}$ of $x_{0}$, there are $G^{F}$-equivariant isomorphisms
$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\mathcal{O}_{{\overline{X}^{S^{*}}(x)}}\Big{)}$
and
$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\omega_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\omega_{{\overline{X}^{S^{*}}(x)}}\Big{)}$
for all $i\geq 0$.
If $C$ is moreover an elliptic conjugacy class in $W$, then the isomorphisms above hold if we replace $x_{0}$ by any other minimal length element in $C$.
Proof.
Fix reduced expressions for $v_{1},\ldots,v_{r}\in W$, and take lifts $v_{1}^{\prime},\ldots,v_{r}^{\prime}\in S^{*}$ corresponding to these reduced expressions. By Theorem 5.1.1, it suffices to show that the cohomology of $X(\underline{v_{1}},\ldots,\underline{v_{r}})$ for the structure sheaf is given by the cohomology of ${\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}$.
Via projection, there is a proper birational morphism $f:{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}\to X(\underline{v_{1}},\ldots,\underline{v_{r}})$ of projective and strongly $F_{p}$-regular schemes (Proposition 2.4.1). Thus there is a $G^{F}$-equivariant isomorphism
$f_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\overset{\sim}{\longrightarrow}\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})},$
by the proof of Zariski’s main theorem in [Hartshorne, Corollary III.11.4].
The definition of local rings being strongly $F_{p}$-regular implies that they are $F_{p}$-rational. Smith has shown that $F_{p}$-rational implies pseudo-rational [Ksmith97, Theorem 3.1]. Moreover, recall that compactified Deligne–Lusztig varieties are normal and Cohen–Macaulay (Proposition 2.4.1). By [Kovacs, Theorem 8.13], the higher direct image $R^{j}f_{*}\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}$ for $j>0$ vanish.
Functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
$H^{i}\Big{(}X(\underline{v_{1}},\ldots,\underline{v_{r}}),\mathcal{O}_{X(\underline{v_{1}},\ldots,\underline{v_{r}})}\Big{)}\overset{\sim}{\longrightarrow}H^{i}\Big{(}{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})},\mathcal{O}_{{\overline{X}^{S^{*}}(v_{1}^{\prime}\cdots v_{r}^{\prime})}}\Big{)},$
for all $i\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from Serre duality for Deligne–Mumford stacks [Nironi, Corollary 2.10].
Remark 5.1.3.
We use the same notations as in Theorem 5.1.1 and set $I={{\text{supp}}}_{F}(x)$. There are $G^{F}$-equivariant morphisms given by [DMR, Proposition 2.3.8]:
${\overline{X}^{S^{*}}(x)}\overset{\sim}{\longrightarrow}G^{F}/U_{I}^{F}\times^{L_{I}^{F}}{\overline{X}^{S^{*}}_{L_{I}}(x)}\overset{\pi_{I}}{\longrightarrow}G^{F}/P_{I}^{F},$
where $P_{I}=BW_{I}B$ is the parabolic subgroup with unipotent radical $U_{I}$ satisfying $P_{I}=L_{I}\ltimes U_{I}$ for a Levi subgroup $L_{I}$ of $G$. Since $G^{F}$ acts on the fibres of $\pi_{I}$ transitively, we have a $G^{F}$-equivariant isomorphism
$H^{i}\Big{(}{\overline{X}^{S^{*}}(x)},\mathcal{O}_{{\overline{X}^{S^{*}}(x)}}\Big{)}\overset{\sim}{\longrightarrow}{\rm Ind}_{P_{I}^{F}}^{G^{F}}H^{i}\Big{(}\overline{X}^{S^{*}}_{L_{I}}(x),\mathcal{O}_{\overline{X}^{S^{*}}_{L_{I}}(x)}\Big{)}.$
for all $i\geq 0$.
5.2. Suzuki and Ree groups
We recall here the assumptions needed on $d$ and $q$ for which $(G,F)$ of type ${}^{2}B_{2}$, ${}^{2}F_{4}$, and ${}^{2}G_{2}$ can arise with $F(S)=S$.
In addition to the assumptions Section 1.1, let $d=2$ and let $q$ be an odd power of $\sqrt{2}$ (resp. $\sqrt{2}$, $\sqrt{3}$). Let $G$ be of type $B_{2}$ (resp. $F_{4}$, $G_{2}$). Then $F:G\to G$ as constructed in Section 1.1 satisfies $F(S)=S$ and $F^{2}$ is the $q^{2}$-Frobenius [DigneMichelBook, §4.3.7]. The group of fixed points $G^{F}$ is the Suzuki group ${}^{2}B_{2}(q)$ (resp. Ree groups ${}^{2}F_{4}(q)$, and ${}^{2}G_{2}(q)$).
Acknowledgements
The author thanks Jochen Heinloth and Georg Linden for their comments on an earlier draft of this paper. Thanks to Ulrich Görtz for helpful conversations.
Parts of the work are carried out when the author was a member of the trimester program "The Arithmetic of the Langlands Program" at the Hausdorff Research Institute for Mathematics in Bonn and when the author was a member of the Research Training Group 2553 in Essen, both funded by the DFG (Deutsche Forschungsgemeinschaft).
[…]
Fakultät für Mathematik, Universität Duisburg-Essen, 45127 Essen, Germany
E-mail address: